### v2

# Domain Adaptation for Offline Reinforcement Learning with Limited Samples

Abstract
Offline reinforcement learning (RL) learns effective policies from a static target dataset. Despite state-of-the-art (SOTA) offline RL algorithms being promising, they highly rely on the quality of the target dataset. The performance of SOTA algorithms can degrade in scenarios with limited samples in the target dataset, which is often the case in real-world applications. To address this issue, domain adaptation that leverages auxiliary samples from related source datasets (such as simulators) can be beneficial. In this context, determining the optimal way to trade off the source and target datasets remains a critical challenge in offline RL. To the best of our knowledge, this paper proposes the first framework that theoretically and experimentally explores how the weight assigned to each dataset affects the performance of offline RL. We establish the performance bounds and convergence neighborhood of our framework, both of which depend on the selection of the weight. Furthermore, we identify the existence of an optimal weight for balancing the two datasets. All theoretical guarantees and optimal weight depend on the quality of the source dataset and the size of the target dataset. Our empirical results on the well-known Procgen Benchmark substantiate our theoretical contributions.
Introduction
Deep reinforcement learning (RL) has demonstrated impressive performance in a wide variety of applications, such as strategy games (Mnih et al. 2013, 2015), robotics (Levine et al. 2016; Duan et al. 2016), and recommender systems (Afsar, Crump, and Far 2022; Lin et al. 2023). RL aims to learn an optimal policy that maximizes the expected cumulative reward. To achieve this goal, the RL agent learns and improves the policy by actively interacting with the environment. However, this poses a critical challenge for the real-world applications of RL, as interactions with the real-world can be significantly dangerous and expensive (Kumar et al. 2020; Levine et al. 2020; Chen, Subramanian, and Paternain 2024). In this context, offline RL merges as a promising alternative framework for the real-world applications of RL where the agent learns effective policies from a static and previously-collected dataset.
Recent advances in offline RL algorithms have shown outstanding performance across a diverse array of problems and datasets (Fujimoto et al. 2019; Kumar et al. 2020; Kostrikov, Nair, and Levine 2021; Chen et al. 2021). Nevertheless, their effectiveness highly depends on the quality of the dataset, such as the amount of data samples. More concretely, it is worth noting that even state-of-the-art (SOTA) offline RL algorithms like BCQ (Fujimoto et al. 2019), CQL (Kumar et al. 2020), IQL (Kostrikov, Nair, and Levine 2021) , DT (Chen et al. 2021) demonstrate poor performance given a small offline RL dataset, as training on a small number of samples may lead to overfitting (Fu et al. 2019; Kumar et al. 2019). In this work, we are interested in offline RL that learns from a static dataset with limited samples.
Figure 1: Schematic of domain adaptation for offline reinforcement learning. The target dataset has limited samples (red), whereas the source dataset has unlimited samples (green) that are with discrepancy (or dynamics gap) from the target domain. Striking the proper balance between the two datasets remains a challenging problem in offline reinforcement learning. The example data in both datasets is the Maze from the Procgen Benchmark (Cobbe et al. 2020).
Related Work
Offline reinforcement learning with dataset distillation.
Dataset distillation (Wang et al. 2018) proposes a framework for synthesizing a smaller and more efficient dataset by minimizing the gradient discrepancy of the samples from the original dataset and the distilled dataset. Synthetic (Light, Liu, and Hu 2024) is the first work that applies dataset distillation to offline RL and achieves comparable performance with the original large offline RL dataset. Specifically, it synthesizes a small distilled dataset by minimizing the gradient matching loss between the original offline RL dataset and the synthetic dataset. However, generating the synthetic dataset necessitates access to the original large offline RL dataset, which is often impractical in real-world scenarios. Indeed, this work focuses on scenarios where only a limited number of samples are accessible.
Offline reinforcement learning with domain adaptation.
To address the overfitting issue by learning from the limited target dataset, domain adaptation techniques (Redko et al. 2020; Farahani et al. 2021) proposes to leverage an auxiliary large source dataset with unlimited samples. H2O (Niu et al. 2022) proposes to improve the learning performance on the limited target dataset, by combining with the unrestricted simulator (the source dataset). However, it is significant to point out that H2O requires knowing the dynamics of both datasets, and the process of training on simulators still requires interacting with the environment online. On the other hand, ORIS (Hou et al. 2024) proposes to generate a new (source) dataset from the simulators, where a generative adversarial network (GAN) model is employed to approximate the state distribution of the original target dataset. Starting from the initial state provided by GAN, the new (source) dataset is generated by interacting with the simulator and reweighted by an additional discriminator model.
Given the unlimited source dataset and the limited target dataset, however, striking the proper balance between the two datasets remains a challenging problem in offline RL (see Figure 1). Solutions involve either combining both datasets equally or using only one of them. For instance, H2O consistently treats both datasets with equal weight. On the other hand, ORIS suggests always assigning a lower weight to the source dataset than to the target dataset, as the dynamics gap could negatively impact the performance of the offline RL. It is worth pointing out that the weights assigned in these cases lack guarantees. To be precise, an equal or lower weight assigned to the source dataset may not achieve optimal performance when the source domain is close to the target domain (in an extreme case they are identical). Consequently, this work focuses on both theoretically and experimentally exploring the impact of the weight of two datasets on the performance of offline RL.
Offline reinforcement learning with dynamics gap.
Dynamics gap in the domain adaptation for offline RL is widely acknowledged as a significant challenge. DARC (Eysenbach et al. 2020) and DARA (Liu, Zhang, and Wang 2022) train a dynamics gap-related penalty by minimizing the divergence between the real and simulator trajectory distributions, and combine it with the simulator reward during the online/offline training of simulators. However, to the best of our knowledge, there have been no prior work that theoretically investigates the impact of the dynamics gap on the performance of offline RL.
Main Contributions
Main contributions of this paper are summarized as follows.
  • To the best of our knowledge, this is the first work that proposes a framework of domain adaptation for offline RL, which allows to both theoretically and experimentally explore the trade-off between the number of samples in a limited target dataset and the discrepancy (or dynamics gap) between the target and source domains.
  • We rigorously establish the performance bounds and convergence neighborhood for our framework. We further identify the existence of an optimal weight for trading off the two datasets, together with the guarantee on the worst-case performance. All theoretical guarantees and the optimal weight will depend on the quality of the source dataset (i.e., dynamics gap) and the number of samples in the target dataset.
  • A series of numerical experiments conducted on the offline Procgen Benchmark (Mediratta et al. 2023) (2D video games) substantiate our theoretical contributions.
Combining Target and Source Datasets
In this section, we consider the mathematical formalism of offline RL (Levine et al. 2020), namely Markov Decision Process (MDP) (Sutton and Barto 2018). MDP is defined by a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},r,\rho,\gamma,T)$, where $\mathcal{S}$ and $\mathcal{A}$ are finite state and action spaces, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to[0,1]$ denotes the transition probability that describes the dynamics of the system, $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ denotes the reward function that evaluates the quality of the decision (action), $\rho:\mathcal{S}\to[0,1]$ denotes the initial state distribution, $\gamma$ represents the discount factor, and $T\in\{0,1,2,\cdots\}$ defines the horizon length.
A policy is represented by the probability distribution over actions conditioned on states. In each state $s$, the agent selects an action $a$ based on the policy $\pi(a|s)$, and generates the next state $s^{\prime}$. The tuple $(s,a,s^{\prime})$ is referred to as the transition data. Offline RL considers employing a behavior policy $\pi_{\beta}$ (or multiple behavior policies) to collect an offline and static dataset with $N$ transitions, i.e., $\hat{\mathcal{D}}=\{(s_{i},a_{i},s_{i}^{\prime})\}_{i=1}^{N},N\in\mathbb{N}_{+}$. The transitions in the dataset $\hat{\mathcal{D}}$ are drawn from a domain $\mathcal{D}$. The goal of RL is to learn an optimal policy $\pi^{*}$ that maximizes the expected (discounted) cumulative reward, i.e.,
$\pi^{*}=\operatorname*{arg\,max}_{\pi}\,\underset{\begin{subarray}{c}s_{0}\sim\rho,\,a_{t}\sim\pi(a_{t}|s_{t})\\ s_{t+1}\sim\mathcal{P}(s_{t+1}|s_{t},a_{t})\end{subarray}}{\mathbb{E}}\left[\sum_{t=0}^{T}\gamma^{t}r(s_{t},a_{t})\right].$ (1)
In the context of offline RL, the policy needs to be learned using exclusively the dataset $\hat{\mathcal{D}}$.
Offline RL algorithms (Fujimoto et al. 2019; Levine et al. 2020; Kostrikov, Nair, and Levine 2021) typically train an action-value function (or Q-function) by minimizing the temporal difference (TD) error iteratively. To be formal, let $\mathcal{B}$ be a Bellman operator. This operator takes different forms depending on the specific algorithm considered. For instance, in Q-learning type methods the Bellman operator takes the form
$\mathcal{B}Q^{k}(s,a)=r(s,a)+\gamma\underset{s^{\prime}\sim\mathcal{P}(s^{\prime}|s,a)}{\mathbb{E}}\left[\max_{a^{\prime}}Q^{k}(s^{\prime},a^{\prime})\right],$ (2)
whereas for actor-critic methods it takes the form
$\mathcal{B}^{\pi}Q^{k}(s,a)=r(s,a)+\gamma\underset{\begin{subarray}{c}s^{\prime}\sim\mathcal{P}(s^{\prime}|s,a)\\ a^{\prime}\sim\pi(a^{\prime}|s^{\prime}),\,\end{subarray}}{\mathbb{E}}\left[Q^{k}(s^{\prime},a^{\prime})\right].$ (3)
Given any transition $(s,a,s^{\prime})$, let us define the TD error with respect to (w.r.t.) this single transition
$\mathcal{E}(Q,(s,a,s^{\prime})):=\left(Q(s,a)-\hat{\mathcal{B}}Q^{k}(s,a)\right)^{2},$ (4)
where $\hat{\mathcal{B}}$ denotes the stochastic approximation of the Bellman operator, namely a version of (2) or (3) where the expectation is replaced by evaluating the random variable at a single of realization $s^{\prime}$. To be precise, the expression on the left-hand side of the above equations depends on the iteration $k$. However, we drop this index to avoid overloading the notation. We further define the expected TD w.r.t. the domain $\mathcal{D}$.
$\mathcal{E}_{\mathcal{D}}(Q)=\underset{(s,a,s^{\prime})\sim\mathcal{D}}{\mathbb{E}}\,\left[\mathcal{E}(Q,(s,a,s^{\prime}))\right].$ (5)
With these definitions, one can define the iterations of a large class of offline RL algorithms through the following optimization problem
$Q^{k+1}=\operatorname*{arg\,min}_{Q}\,\mathcal{E}_{\mathcal{D}}(Q).$ (6)
It is worth pointing out that the specific forms of (2) and (3) can result in poor performance in offline RL attributed to the issues with bootstrapping from out-of-distribution (OOD) actions (Fujimoto, Meger, and Precup 2019; Kumar et al. 2019, 2020; Levine et al. 2020). This typically leads to an overestimation of the Q-value. To avoid this overestimation, prior works consider solely using in-distribution state-action pairs to maintain the Q-function (Fujimoto et al. 2019), or constraining the learned policy to remain closely aligned with the behavior policy (Levine et al. 2020). These algorithms, however, can also be written as (5) by introducing modifications to the Bellman operators.
Note that the expectation in (5) poses a challenge in solving problem (6): it requires visiting every transition infinite times. In practice, one can only work with the empirical version of the TD error $\mathcal{E}_{\mathcal{D}}(Q)$ in (5), i.e.,
$\mathcal{E}_{\hat{\mathcal{D}}}(Q)=\frac{1}{N}\sum_{i=1}^{N}\mathcal{E}(Q,(s_{i},a_{i},s_{i}^{\prime})),$ (7)
where the samples are from dataset $\hat{\mathcal{D}}$. Then, the offline RL algorithm is defined as the minimization of the stochastic approximation of problem (6)
$\hat{Q}^{k+1}=\operatorname*{arg\,min}_{Q}\,\mathcal{E}_{\hat{\mathcal{D}}}(Q).$ (8)
It has been widely demonstrated that state-of-the-art offline RL algorithms, such as BCQ (Fujimoto et al. 2019), CQL (Kumar et al. 2020), IQL (Kostrikov, Nair, and Levine 2021), and DT (Chen et al. 2021), are capable of solving problem (6) given sufficient transition samples from the domain $\mathcal{D}$. Nevertheless, this assumption may not be realizable in practice, e.g., healthcare (Tang et al. 2022) and autonomous driving (Pan et al. 2017), where data-collection is challenging. In this regime, $\hat{Q}^{k+1}$ generally demonstrates poor performance in approximating $Q^{k+1}$, as training on a small number of samples may lead to overfitting (Fu et al. 2019; Kumar et al. 2019).
On the other hand, in some applications, one can rely on simulators (or related datasets) that provide a larger number of samples $\mathcal{D}^{\prime}=\{(s_{j},a_{j},s_{j}^{\prime})\}_{j=1}^{N^{\prime}}$ with $N^{\prime}\gg N$. It is worth noting that, in general, $\mathcal{D}^{\prime}$ will differ from $\mathcal{D}$ in terms of the state distribution and transition probabilities. We term $\mathcal{D}^{\prime}$, $\hat{\mathcal{D}}$ and $\mathcal{D}$ the source dataset, the target dataset, and the target domain respectively. Similar to (5), we define
$\mathcal{E}_{\mathcal{D}^{\prime}}(Q)=\underset{(s_{j},a_{j},s_{j}^{\prime})\sim\mathcal{D}^{\prime}}{\mathbb{E}}\,\left[\mathcal{E}(Q,(s_{j},a_{j},s_{j}^{\prime}))\right].$ (9)
To be precise, one still needs to work with the empirical TD error on the dataset $\mathcal{D}^{\prime}$ since estimating (9) is infeasible. Note however, that since $N^{\prime}\gg N$ the generalization error due to having a finite number of samples in the source dataset $\mathcal{D}^{\prime}$ is negligible. For this reason and to simplify the exposition we choose to assume that one can compute (9), and denote by $\mathcal{D}^{\prime}$ both the source dataset and the source domain with a slight abuse of notation.
Prior works such as H2O (Niu et al. 2022) and ORIS (Hou et al. 2024) propose to improve the performance of the limited target dataset, by combining with a far larger source dataset. Significantly, H2O consistently treats both datasets equally, whereas ORIS always assigns a lower weight to the source dataset. In this paper, we explore a more general approach to combine both datasets, i.e.,
$Q_{\lambda}^{k+1}=\operatorname*{arg\,min}_{Q}\,(1-\lambda)\mathcal{E}_{\hat{\mathcal{D}}}(Q)\!+\!\lambda\mathcal{E}_{\mathcal{D}^{\prime}}(Q),$ (10)
where $\lambda\in[0,1]$ denotes the weight that trades off $\mathcal{E}_{\hat{\mathcal{D}}}(Q)$ from the target dataset and $\mathcal{E}_{\mathcal{D}^{\prime}}(Q)$ from the source dataset. Particularly, (10) recovers H2O when $\lambda=0.5$ and yields ORIS when $\lambda<0.5$. Indeed, $\lambda\approx 1$ emphasizes the minimization of TD error corresponding in the source dataset. This approach is suitable in cases where the source dataset is similar to the target (or coming from the same domain in an extreme case). On the other hand, $\lambda\approx 0$ focuses on minimizing the TD error in the target dataset. This approach is suitable in cases where data is abundant (large $N$) or when the discrepancies between the target and source datasets are too large.
Given these observations, it is expected that different values of $\lambda$ attain optimal performance depending on the interplay between the available number of target samples and the discrepancies between the datasets. The following section formalizes this claim.
Performance and Convergence Guarantees
We start this section by discussing the necessary assumptions to develop our theoretical results concerning the generalization of the solution to problem (10).
Assumption 1.
The absolute value of the reward $r(s,a)$ is bounded by a constant $B>0$ for all state-action pairs.
Assumption 1 is common in the literature (Azar, Osband, and Munos 2017; Wei et al. 2020; Zhang, Du, and Ji 2021).In particular, in the case of finite state-action spaces, it is always possible to design the reward to avoid the possibility of being unbounded. Note that this assumption along with the fact that a geometric series with ratio $\gamma\in[0,1)$ converges to $1/(1-\gamma)$ implies that
$\max_{(s,a)\in\mathcal{S}\times\mathcal{A}}Q(s,a)\leq\frac{B}{1-\gamma}.$ (11)
To proceed, we denote by $P_{\mathcal{D^{\prime}}}$ and $P_{\mathcal{D}}$ the probability of a state-action pair $(s,a)$ from the source and the target domains. Define $N(s,a)\in\left\{0,1,\ldots,N\right\}$ to be the number of samples in the target dataset for the state-action pair $(s,a)$.
Assumption 2.
The source dataset, the target dataset, and the target domain have the same state-action distribution. This is, $\forall(s,a)\in\mathcal{S}\times\mathcal{A}$, there exists a $P(s,a)>0$ such that
$P(s,a)=P_{\mathcal{D^{\prime}}}(s,a)=P_{\mathcal{D}}(s,a)=\frac{N(s,a)}{N}.$
Assumption 2 contains two assumptions. First, it assumes that the probability of sampling a given state and action in both domains is equal. This assumption can be met in cases where the data in $\mathcal{D}^{\prime}$ is generated with a simulator, as we have the freedom to choose the starting point for any transition. Note that, the transition probabilities in the two domains are still different, i.e., $P_{\mathcal{D}}(s^{\prime}|s,a)\neq P_{\mathcal{D^{\prime}}}(s^{\prime}|s,a)$. This is referred to as dynamics gap (Niu et al. 2022; Hou et al. 2024). Additionally, we also assume that the ratio of samples for a specific state-action pair over the total number of samples is equal to the probability of sampling that specific state-action. This assumption is a good approximation as $N$ goes to infinity. We chose to keep this assumption for the simplicity of the exposition as its only consequence is to simplify factors $P(s,a)N/N(s,a)$ to one.
Finally, note that the transition sample $\tau$ are independent and identically distributed (i.i.d.), as the dataset $\hat{\mathcal{D}}$ is shuffled and the transition data are sampled i.i.d.. Alternatively, one can think of the dataset as random samples from the occupancy measure, followed by a random transition.
We next define two quantities that will play an important role in the bounds that we establish. Start by defining
$\xi=\underset{(s,a)\in\mathcal{S}\times\mathcal{A}}{\max}\left[\left(\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)-\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)\right)^{2}\right].$ (12)
The quantity at hand measures the discrepancy or dynamics gap between the two domains. Indeed, if the transition probabilities in the source and target domains match, the quantity in (12) is zero. We also define a measure of the variability of the target dataset. Let $\sigma^{2}\left(\hat{\mathcal{B}}Q^{k}(s,a)\right)$ represent the variance of the stochastic Bellman operator for the state-action pair $(s,a)$. Then we define the following quantity
$\varsigma=\underset{(s,a)\in\mathcal{S}\times\mathcal{A}}{\max}\frac{\sigma^{2}\left(\hat{\mathcal{B}}Q^{k}(s,a)\right)}{N(s,a)}.$ (13)
The term above is the maximum normalized variance across all state-action pairs. Indeed, note that the number of samples to keep the value $\varsigma$ constant is proportional to the variance.
With these assumptions and definitions, we are in conditions of establishing the first result of this work which bounds the expected value of the TD error for $Q_{\lambda}^{k+1}$, defined in (10), on the target domain.
Theorem 1 (Expected Performance Bound).
Let Assumption 1 and Assumption 2 hold. Recall $\xi$ in (12) and $\varsigma$ in (13), and that $Q^{k+1}$ and $Q_{\lambda}^{k+1}$ denote the solutions to (6) and (10). For all iterations $(k=0,1,2,\cdots)$, it holds that
$\mathbb{E}\left[\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})\right]-\mathcal{E}_{\mathcal{D}}(Q^{k+1})\leq(1-\lambda)^{2}\varsigma+\lambda^{2}\xi.$ (14)
The previous theorem establishes the bias-variance trade-off sought by combining the two datasets with different weights in (10). Indeed, note that the optimal weight that minimizes the right-hand-side of the above bound is
$\lambda^{*}=\frac{\varsigma}{\varsigma+\xi}.$ (15)
This recovers the intuition that if the dataset has no variation (or the number of samples is large) $\varsigma\approx 0$, which implies that in (10) only the target dataset is considered. On the other hand, when the variance on the target dataset is large ($\varsigma\to\infty$) or if the two domains are similar ($\xi\approx 0$), the optimal value of $\lambda$ is one. This suggests that one should use the source domain. Although intuitive, the above bound is insufficient to claim any generalization guarantees as the tails of the distribution could be heavy. The next theorem provides the generalization bound.
Theorem 2 (Worst-Case Performance Bound).
Under the hypotheses of Theorem 1, the following bound holds with probability $1-\delta$
$\mathcal{E}_{\mathcal{D}}\left(Q_{\lambda}^{k+1}\right)-\mathcal{E}_{\mathcal{D}}\left(Q^{k+1}\right)\leq(1-\lambda)^{2}\varsigma+\lambda^{2}\xi$ (16)
$+\sqrt{\frac{\log\left(1/\delta\right)}{2}}\frac{|\mathcal{S}||\mathcal{A}|}{\sqrt{N}}\frac{\gamma B(1-\lambda)}{(1-\gamma)}\left(8\frac{B(1-\lambda)}{(1-\gamma)}+4\lambda\sqrt{\xi}\right).$
In particular, when $\lambda$ is selected as in (15), this bound takes the following form
$\mathcal{E}_{\mathcal{D}}\left(Q_{\lambda^{*}}^{k+1}\right)-\mathcal{E}_{\mathcal{D}}\left(Q^{k+1}\right)\leq\frac{\varsigma\xi}{\varsigma+\xi}+$ (17)
$\sqrt{\frac{\log\left(1/\delta\right)}{2}}\frac{|\mathcal{S}||\mathcal{A}|}{\sqrt{N}}\frac{\gamma B}{(1-\gamma)}\frac{\xi}{\varsigma+\xi}\!\!\left(\!\frac{8B}{(1-\gamma)}\frac{\xi}{\varsigma+\xi}\!+\!4\frac{\varsigma\sqrt{\xi}}{\varsigma+\xi}\right).$
The previous theorem provides the worst-case bound of solving (10), which demonstrates the bias-variance trade-off by the two datasets with different weights as well. We also establish in (17) the worst-case bound of applying $\lambda^{*}$ in (15).
(a) Caveflyer
(b) Climber
(c) Dodgeball
(d) Maze
(e) Miner
Figure 2: The performance of offline RL with domain adaptation across five Procgen games. The source dataset contains $40000$ samples from levels $[0,99]$, while the target dataset comprises $1000$ samples from levels $[100,199]$. We consider seven weights, $\lambda\in\{0,0.2,0.4,0.5,0.6,0.8,1.0\}$, to trade off the source and target datasets with the star marking the optimal weight.
Akin to (15), the optimal weight that minimizes the right-hand-side of the bound (16) in Theorem 2 is
$\hat{\lambda}^{*}=\frac{\varsigma+\sqrt{\frac{\log\left(1/\delta\right)}{2}}\frac{|\mathcal{S}||\mathcal{A}|}{\sqrt{N}}\frac{\gamma B}{1-\gamma}\left(\frac{8B}{1-\gamma}-2\sqrt{\xi}\right)}{\varsigma+\xi+\sqrt{\frac{\log\left(1/\delta\right)}{2}}\frac{|\mathcal{S}||\mathcal{A}|}{\sqrt{N}}\frac{\gamma B}{1-\gamma}\left(\frac{8B}{1-\gamma}-4\sqrt{\xi}\right)}.$ (18)
It is worth pointing out that when $\xi=0$, (18) simplifies to $\hat{\lambda}^{*}=1$, consistent with (15). However, when $\varsigma=0$, $\hat{\lambda}^{*}$ is not necessarily to be 0. This suggests that the worst-case bound may be conservative.
More importantly, both optimal weights for the expected performance bound and the worst-case performance bound, as shown in (15) and (18), implies that the optimal trade-off between the source and target datasets is not always trivial, meaning that $\lambda^{*}$ may not belong to $\{0,0.5,1\}$. The optimal trade-off for the expected performance depends on the number of samples in the target dataset (corresponding to $\varsigma$) and the discrepancy (or dynamics gap) between the two domains (corresponding to $\xi$). In addition to $\varsigma$ and $\xi$, the optimal weight for the worst-case performance will depend on more factors such as the reward bound $B$, the size of the state and action spaces $|\mathcal{S}|$ and $|\mathcal{A}|$.
Having established the performance bounds of solving (10), we are in conditions of providing the convergence guarantee. We formalize it in the next theorem, which relies on the following two quantities: maximum of dynamics gap as in (12) and maximum of standard deviation as in (13).
$\xi_{\text{max}}=\sup_{k\in\mathbb{N}}\,\xi(Q^{k}),\leavevmode\nobreak\ \sigma_{\text{max}}=\sup_{k\in\mathbb{N}}\,\underset{(s,a)\in\mathcal{S}\times\mathcal{A}}{\max}\sigma\left(\hat{\mathcal{B}}Q^{k}(s,a)\right).$ (19)
Theorem 3 (Convergence).
Under the hypotheses of Theorem 1, it holds that
[…]
$\leq\gamma^{k+1}\!\!\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[||Q^{0}(s,a)-Q^{*}(s,a)||_{\infty}\right]$
$+\frac{1-\gamma^{k+1}}{1-\gamma}\left((1-\lambda)\sigma_{\text{max}}+\lambda\sqrt{\xi_{\text{max}}}\right).$ (20)
The previous theorem implies that the expected solution of solving (10) is guaranteed to converge to a neighborhood of the optimal $Q$-function, i.e., $Q^{*}$ as $k\to\infty$. This neighborhood is presented as follows
$\mathcal{C}=\frac{(1-\lambda)\sigma_{\text{max}}+\lambda\sqrt{\xi_{\text{max}}}}{1-\gamma}.$ (21)
In addition to the discount factor $\gamma$, the neighborhood $\mathcal{C}$ depends on the weight $\lambda$, the maximal dynamics gap $\xi_{\text{max}}$, and the maximal standard deviation $\sigma_{\text{max}}$. The next section presents a series of numerical experiments to validate the corresponding theoretical contributions in this section.
(a) Caveflyer
(b) Climber
(c) Dodgeball
(d) Maze
(e) Miner
Figure 3: The performance of offline RL with domain adaptation across five Procgen games. The target dataset comprises $1000$ samples from levels $[100,199]$, and three source datasets are considered, each containing $40000$ samples from levels $[0,99]$ (green, $\xi_{1}$), $[25,124]$ (blue, $\xi_{2}$), and $[50,149]$ (red, $\xi_{3}$), respectively. Seven weights, $\lambda\in\{0,0.2,0.4,0.5,0.6,0.8,1.0\}$, are evaluated to trade off the source and target datasets with the star marking the optimal weight for each $\xi$.
(a) Caveflyer
(b) Climber
(c) Dodgeball
(d) Maze
(e) Miner
Figure 4: The performance of offline RL with domain adaptation across five Procgen games. The source dataset comprises $40000$ samples from levels $[50,149]$, and target datasets from levels $[100,199]$ are considered with three different sample sizes: $N=1000$ (green), $N=2500$ (blue), and $N=4000$ (red). Seven weights, $\lambda\in\{0,0.2,0.4,0.5,0.6,0.8,1.0\}$, are evaluated to trade off the source and target datasets with the star marking the optimal weight for each $N$.
Numerical Experiments
In this section, we demonstrate the numerical performance of solving (10) under different weight $\lambda$, as well as validating the theoretical contributions in the previous section.
Environmental Setup
We consider the well-known offline Procgen Benchmark (Mediratta et al. 2023) that is often used to assess the domain adaptation/generalization of offline RL. We select five games from Procgen: Caveflyer, Climber, Dodgeball, Maze, Miner. The environmental details including the introduction of the five games are deferred to the appendix.
As discussed in the Introduction, our domain adaptation framework for offline RL explores the trade-off between the limited target dataset and the unlimited source dataset. Consequently, this framework is naturally applicable to various state-of-the-art offline RL algorithms, such as BCQ (Fujimoto et al. 2019), CQL (Kumar et al. 2020), IQL (Kostrikov, Nair, and Levine 2021), and DT (Chen et al. 2021). In this work, we select CQL (Kumar et al. 2020) as a representative algorithm, as it generally demonstrates a promising and robust performance across a variety of offline RL problems and tasks (Kumar et al. 2020; Fu et al. 2020). To ensure a fair comparison, we retain consistent all hyperparameters, e.g., batch size, learning rate, and network size, and solely change the weight w.r.t. each of the datasets. The implementation details are provided in the appendix.
Results
Recall that the worst-case performance bound in (16) can be considerably conservative, particularly when the size of the state and action spaces are very large. Therefore, this subsection focuses on the expected performance bound (14) as well as its corresponding corollary (15). It is crucial to note that the variance $\varsigma$ and the dynamics gap $\xi$ in (14) are challenging to measure or estimate precisely, as it requires access to the entire source and target domains. This is not possible in our problem of interest. Nevertheless, one can still investigate how these factors influence the expected bound (14), which provides insight into the expected performance of offline RL. Specifically, we examine the impact of each of $\varsigma$, $\xi$, and $\lambda$ on the expected performance. To achieve this, we vary one of these factors at a time while keeping the other two parameters constant.
Note that Procgen employs procedural content generation to create adaptive levels upon episode reset. Each level corresponds to a specific seed (non-negative integer) and has distinct dynamics and visual appearances. We consider the source and target domains to be with levels $[0,99]$ and $[100,199]$, respectively. Recall that the target dataset is expected to contain significantly fewer samples than the source dataset, i.e., $N\ll N^{\prime}$. Typically, $N^{\prime}$ is considered to be ten times larger than $N$. Therefore, we set $N=1000$ and sample from levels $[100,199]$ for the target dataset, and set $N^{\prime}=40000$ and sample from levels $[0,99]$ for the source dataset. We denote by $\xi_{1}$ the discrepancy (or dynamics gap) between levels $[0,99]$ and $[100,199]$
Impact of $\lambda$.
Table 1: The performance across all $\lambda\in\{0,0.2,0.4,0.5,0.6,0.8,1.0\}$ (mean and standard deviation) with fixed $N=1000$ corresponding to different $\xi$.
Game $\xi_{1}([0,99])$ $\xi_{2}([25,124])$ $\xi_{3}([50,149])$
Caveflyer $1.23\pm 0.23$ $1.28\pm 0.26$ $\mathbf{1.37\pm 0.34}$
Climber $0.57\pm 0.10$ $0.73\pm 0.16$ $\mathbf{0.85\pm 0.21}$
Dodgeball $0.67\pm 0.17$ $0.87\pm 0.21$ $\mathbf{1.29\pm 0.42}$
Maze $3.52\pm 0.41$ $3.97\pm 0.42$ $\mathbf{4.35\pm 0.52}$
Miner $1.27\pm 0.48$ $2.05\pm 0.55$ $\mathbf{3.12\pm 0.80}$
We consider seven $\lambda$ from $\{0,0.2,0.4,0.5,0.6,0.8,1\}$, where a larger $\lambda$ focuses more on the source dataset, as implied by (10). Note that $\lambda=0$ trains on the limited target dataset only, whereas $\lambda=1$ solely considers the unlimited source dataset. $\lambda=0.5$ represents the equal combination of the two dataset. These three are the trivial choices of $\lambda$ in general. (14) reveals that the optimal weight that minimizes the expected performance bound is not necessarily to be 0, or 0.5 or 1. Indeed, our experiments on the Procgen, as depicted in Figure 2, validates this claim by the observation that only one out of five games has the optimal weight to be the trivial choice. In addition, Miner shows that all three trivial choices of $\lambda$ achieve worse performance compared to other weights.
Impact of $\xi$.
We consider three source datasets of the same size but from different domains: levels $[0,99]$, levels $[25,124]$, $[50,149]$. Similar to $\xi_{1}$, let $\xi_{2}$ and $\xi_{3}$ represent the discrepancies (or dynamics gap) between levels $[25,124]$ and $[100,199]$, and between levels $[50,149]$ and $[100,199]$, respectively. Thus, we obtain $\xi_{1}\geq\xi_{2}\geq\xi_{3}$, as more overlap between the levels of the two datasets demonstrates smaller discrepancies between them. The bound (14) in Theorem 1 implies that the expected performance improves as $\xi$ decreases. The numerical results of five games across three different $\xi$ values are summarized in Table 1, which supports the implication from the bound (14).
Impact of $\varsigma$.
Given $\varsigma$ in (13), it is intuitive that the normalized variance $\varsigma$ decreases with $N(s,a)$. Because of the positive proportional relationship between $N$ and $N(s,a)$, a larger $N$ is expected to lead to a smaller $\varsigma$. The bound (14) in Theorem 1 implies that the expected performance improves as $\varsigma$ decreases. The numerical results of five games across three different $N$ values are summarized in Table 2, which validates the implication from the bound (14).
Table 2: The performance across all $\lambda\in\{0,0.2,0.4,0.5,0.6,0.8,1.0\}$ (mean and standard deviation) with fixed $\xi_{3}$ corresponding to different $N$.
Game $N=1000$ $N=2500$ $N=4000$
Caveflyer $1.37\pm 0.34$ $1.52\pm 0.26$ $\mathbf{1.57\pm 0.19}$
Climber $0.85\pm 0.21$ $0.89\pm 0.29$ $\mathbf{0.95\pm 0.26}$
Dodgeball $1.29\pm 0.42$ $1.34\pm 0.30$ $\mathbf{1.42\pm 0.18}$
Maze $4.35\pm 0.52$ $4.75\pm 0.59$ $\mathbf{5.21\pm 0.69}$
Miner $3.12\pm 0.80$ $3.28\pm 0.59$ $\mathbf{3.33\pm 0.75}$
Optimal Weight $\lambda^{*}$.
$\lambda^{*}$ in (15) implies that the optimal value of $\lambda$ equals to zero when the dataset has no variation (or the number of samples $N$ is very large) $\varsigma\approx 0$. This suggests to use the target dataset only. On the other hand, if the variance on the target dataset is very large ($\varsigma\to\infty$) or if the two domains are similar ($\xi\approx 0$), the optimal weight simplifies to one, meaning that one should consider the source dataset only. Although intuitive, it is important to note that (15) represents the optimal solution for minimizing the expected performance bound, instead of the expected performance itself. Thus, there is no guarantee that $\lambda^{*}=1$ only if $\xi=0$ or $\varsigma\to\infty$, nor that $\lambda^{*}=0$ only if $\varsigma=0$ or $\xi\to\infty$. Moreover, this subsection seeks to identify the optimal value of $\lambda$ from the discrete set $\lambda\in\{0,0.2,0.4,0.5,0.6,0.8,1\}$.
$\lambda^{*}$ varies with $\xi$.
Table 3: The optimal weight corresponding to different $\xi$.
Game $\xi_{1}([0,99])$ $\xi_{2}([25,124])$ $\xi_{3}([50,149])$
Caveflyer $0.4$ $0.4$ $0.5$
Climber $0.6$ $0.8$ $1.0$
Dodgeball $0.5$ $0.6$ $0.8$
Maze $0.4$ $0.6$ $0.8$
Miner $0.4$ $0.6$ $0.8$
To explore how $\lambda^{*}$ varies with $\xi$, we fix $N=1000$ in each of the five Procgen games and select three different source datasets with $\xi_{1}\geq\xi_{2}\geq\xi_{3}$. We then implement this using seven values of $\lambda\in\{0,0.2,0.4,0.5,0.6,0.8,1\}$. The numerical results are summarized in Figure 3 and Table 3. Both results demonstrate that the optimal weight $\lambda^{*}$ within $\{0,0.2,0.4,0.5,0.6,0.8,1\}$ increases as $\xi$ decreases. This substantiates the intuition from Theorem 1 and (15) that greater emphasis should be placed on the source dataset when its discrepancy from the target domain is smaller. In the extreme case, where the source and target domains are identical, one should consider the source dataset only.
$\lambda^{*}$ varies with $\varsigma$.
Table 4: The optimal weight corresponding to different $N$.
Game $N=1000$ $N=2500$ $N=4000$
Caveflyer $0.5$ $0.5$ $0.2$
Climber $1.0$ $0.8$ $0.4$
Dodgeball $0.8$ $0.6$ $0.5$
Maze $0.8$ $0.4$ $0.2$
Miner $0.8$ $0.8$ $0.6$
To explore how $\lambda^{*}$ varies with $\varsigma$, we fix $\xi_{3}$ in each of the five Procgen games. Since a larger $N$ corresponds to a smaller $\varsigma$, we select three different values of $N\in\{1000,2500,4000\}$. It is worth noting that $N^{\prime}$ is consistently maintained at ten times larger than $N$, ensuring that the source dataset always contains significantly more samples than the target dataset. We then implement this using seven values of $\lambda\in\{0,0.2,0.4,0.5,0.6,0.8,1\}$. The numerical results are summarized in Figure 4 and Table 4. Both results demonstrate that the optimal weight $\lambda^{*}$ within $\{0,0.2,0.4,0.5,0.6,0.8,1\}$ decreases as $N$ increases (i.e., $\varsigma$ decreases). This substantiates the intuition from Theorem 1 and (15) that greater emphasis should be placed on the target dataset when it contains more samples. In the extreme case, where the target dataset has an infinite number of samples, one should solely use the target dataset.
Conclusion
Offline RL is highly dependent on the size of the target dataset. Even state-of-the-art offline RL algorithms often lack performance guarantees under a limited number of samples. To tackle offline RL with limited samples, domain adaptation can be employed which considers related source datasets e.g., simulators that typically offer unlimited (or a very large number of) samples. To the best of our knowledge, we propose in this paper the first framework that theoretically and experimentally explores the domain adaptation for offline RL with limited samples. Specifically, we establish the expected and worst-case performance bounds, as well as a convergence neighborhood under our framework. Moreover, this work provides the optimal weight for trading off the unlimited source dataset and the limited target dataset. It demonstrates that the optimal weight is not necessarily one of the trivial choices: using either dataset or combining the two datasets equally. We conduct a series of numerical experiments on the renowned benchmark Procgen (2D video games). Our numerical results validate the theoretical contributions in this work.
Acknowledgments
This work was sponsored by the Office of Naval Research (ONR), under contract number N00014-23-1-2377.
Appendix A Appendix
In this appendix, we provide the proofs of our theoretical contributions together with experimental details in this work. We start by proving Theorem 1. To proceed, we rely on the following proposition.
Proposition 1.
Recall the empirical Bellman operator $\hat{\mathcal{B}}$ in (4). Denote by $\mathcal{B}_{\mathcal{D}}$ ($\mathcal{B}_{\mathcal{D^{\prime}}}$) the Bellman operator in (2) or (3) in which $s^{\prime}$ follows the transition probability of the domain $\mathcal{D}$ ($\mathcal{D^{\prime}}$). Recall $N(s,a)$ in Assumption 2, and that $Q^{k+1}$ and $Q_{\lambda}^{k+1}$ represent the solutions to (6) and (10). At each iteration $(k=1,2,3,\cdots)$, it holds that
$Q^{k+1}(s,a)=\mathcal{B}_{\mathcal{D}}Q^{k}(s,a),\,\forall(s,a)\in\mathcal{D},$ (22)
$Q_{\lambda}^{k+1}(s,a)=\frac{1-\lambda}{N(s,a)}\sum_{j=1}^{N(s,a)}\hat{\mathcal{B}}_{s_{j}^{\prime}}Q^{k}(s,a)+\lambda\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a),\,\forall(s,a)\in\mathcal{D}.$ (23)
Proof.
We start by proving (22). For any $Q$, we note that
$\mathcal{E}_{\mathcal{D}}(Q)$ =$\stackrel{{\scriptstyle(a)}}{{=}}\underset{(s,a,s^{\prime})\sim\mathcal{D}}{\mathbb{E}}\left[\mathcal{E}(Q,(s,a,s^{\prime}))\right]$ (24)
=$\stackrel{{\scriptstyle(b)}}{{=}}\underset{(s,a,s^{\prime})\sim\mathcal{D}}{\mathbb{E}}\left[\left(Q(s,a)-\hat{\mathcal{B}}Q^{k}(s,a)\right)^{2}\right]$ (25)
=$\stackrel{{\scriptstyle(c)}}{{=}}\sum_{(s,a,s^{\prime})\sim\mathcal{D}}P_{\mathcal{D}}(s,a,s^{\prime})\left(Q(s,a)-\hat{\mathcal{B}}Q^{k}(s,a)\right)^{2}$ (26)
where $(a)$ is given by (5), $(b)$ is by definition in (4), $(c)$ follows from the definition of expectation.
The derivative of $\mathcal{E}_{\mathcal{D}}(Q)$ w.r.t. a single $Q(s,a),\forall(s,a)\in\mathcal{D}$ is given by
$\frac{\partial\mathcal{E}_{\mathcal{D}}(Q)}{\partial Q(s,a)}$ =$\stackrel{{\scriptstyle(a)}}{{=}}\sum_{s^{\prime}}P_{\mathcal{D}}(s,a,s^{\prime})2\left(Q(s,a)-\hat{\mathcal{B}}Q^{k}(s,a)\right)$ (27)
=$\stackrel{{\scriptstyle(b)}}{{=}}2\sum_{s^{\prime}}P(s,a)P_{\mathcal{D}}(s^{\prime}|s,a)\left(Q(s,a)-\hat{\mathcal{B}}Q^{k}(s,a)\right)$ (28)
=$\stackrel{{\scriptstyle(c)}}{{=}}2P(s,a)\sum_{s^{\prime}}P_{\mathcal{D}}(s^{\prime}|s,a)\left(Q(s,a)-\hat{\mathcal{B}}Q^{k}(s,a)\right)$ (29)
=$\stackrel{{\scriptstyle(d)}}{{=}}2P(s,a)\left(\sum_{s^{\prime}}P_{\mathcal{D}}(s^{\prime}|s,a)Q(s,a)-\sum_{s^{\prime}}P_{\mathcal{D}}(s^{\prime}|s,a)\hat{\mathcal{B}}Q^{k}(s,a)\right)$ (30)
=$\stackrel{{\scriptstyle(e)}}{{=}}2P(s,a)\left(Q(s,a)-\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)\right).$ (31)
where $(a)$ follows from algebraic manipulations, $(b)$ is by splitting the joint probability $P_{\mathcal{D}}(s,a,s^{\prime})$ and Assumption 2, $(c)$ is due to that $P(s,a)$ is independent of $s^{\prime}$, $(d)$ follows from algebraic manipulations, $(e)$ is due to that $Q(s,a)$ is independent of $s^{\prime}$, the probability sums to 1, and the definition $\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)=\mathbb{E}_{s^{\prime}\sim\mathcal{P}_{\mathcal{D}}(s^{\prime}|s,a)}\left[\hat{\mathcal{B}}Q^{k}(s,a)\right]$.
Since $Q^{k+1}$ is the solution to (6), we obtain
$\frac{\partial\mathcal{E}_{\mathcal{D}}(Q^{k+1})}{\partial Q^{k+1}(s,a)}=0,\,\forall(s,a)\in\mathcal{D}.$ (32)
Combining the previous equation with (31) and the fact that $P(s,a)>0$ (by Assumption 2) yields
$Q^{k+1}(s,a)=\mathcal{B}_{\mathcal{D}}Q^{k}(s,a),\,\forall(s,a)\in\mathcal{D}.$ (33)
This completes the proof of (22) in Proposition 1.
We now turn to derive $Q_{\lambda}^{k+1}(s,a),\forall(s,a)\in\mathcal{D}$. Let us define
$F$ $=(1-\lambda)\mathcal{E}_{\hat{\mathcal{D}}}(Q)+\lambda\mathcal{E}_{\mathcal{D}^{\prime}}(Q)$ (34)
$=(1-\lambda)\frac{1}{N}\sum_{i=1}^{N}\left(Q(s_{i},a_{i})-\hat{\mathcal{B}}Q^{k}(s_{i},a_{i})\right)^{2}+\lambda\mathcal{E}_{\mathcal{D}^{\prime}}(Q)$ (35)
where the previous equation follows from (4) and (7).
The derivative of $F$ w.r.t. a single $Q(s,a),\forall(s,a)\in\mathcal{D}$ is given by
$\frac{\partial F}{\partial Q(s,a)}$ =$\stackrel{{\scriptstyle(a)}}{{=}}(1-\lambda)\frac{2}{N}\sum_{j=1}^{N(s,a)}\left(Q(s,a)-\hat{\mathcal{B}}_{s_{j}^{\prime}}Q^{k}(s,a)\right)+2\lambda P_{\mathcal{D^{\prime}}}(s,a)\left(Q(s,a)-\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)\right)$ (36)
=$\stackrel{{\scriptstyle(b)}}{{=}}(1-\lambda)\frac{2}{N}\left(N(s,a)Q(s,a)-\sum_{j=1}^{N(s,a)}\hat{\mathcal{B}}_{s_{j}^{\prime}}Q^{k}(s,a)\right)+2\lambda P(s,a)\left(Q(s,a)-\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)\right)$ (37)
=$\stackrel{{\scriptstyle(c)}}{{=}}2(1-\lambda)\left(P(s,a)Q(s,a)-\frac{P(s,a)}{N(s,a)}\sum_{j=1}^{N(s,a)}\hat{\mathcal{B}}_{s_{j}^{\prime}}Q^{k}(s,a)\right)+2\lambda P(s,a)\left(Q(s,a)-\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)\right)$ (38)
where the first term in $(a)$ follows from algebraic manipulations and the fact that only the quantities w.r.t. $(s,a)$-pair survive, and the second term in $(a)$ follows from the derivation of $\partial\mathcal{E}_{\mathcal{D}}(Q)/\partial Q(s,a)$ solely replacing $\mathcal{D}$ by $\mathcal{D^{\prime}}$, (b) is due to that $Q(s,a)$ is independent of $s^{\prime}$ and the fact that $P_{\mathcal{D^{\prime}}}(s,a)=P(s,a)$ (by Assumption 2), (c) directly follows from Assumption 2, which indicates that $N(s,a)/N=P(s,a)$.
Since $Q_{\lambda}^{k+1}$ is the solution to (10) and re-arranging (38), we obtain
$Q_{\lambda}^{k+1}(s,a)\left((1-\lambda)P(s,a)+\lambda P(s,a)\right)=(1-\lambda)\frac{P(s,a)}{N(s,a)}\sum_{j=1}^{N(s,a)}\hat{\mathcal{B}}_{s_{j}^{\prime}}Q^{k}(s,a)+\lambda P(s,a)\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a).$ (39)
Since $P(s,a)>0$, we further have
$Q_{\lambda}^{k+1}(s,a)=\frac{1-\lambda}{N(s,a)}\sum_{j=1}^{N(s,a)}\hat{\mathcal{B}}_{s_{j}^{\prime}}Q^{k}(s,a)+\lambda\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a),\,\forall(s,a)\in\mathcal{D}.$ (40)
This completes the proof of Proposition 1. ∎
Having introduced Proposition 1, we are in conditions of proving Theorem 1.
Proof of Theorem 1
Theorem 1 (Expected Performance Bound). Let Assumption 1 and Assumption 2 hold. Recall $\xi$ in (12) and $\varsigma$ in (13), and that $Q^{k+1}$ and $Q_{\lambda}^{k+1}$ denote the solutions to (6) and (10). For all iterations $(k=1,2,3,\cdots)$, it holds that
$\mathbb{E}\left[\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})\right]-\mathcal{E}_{\mathcal{D}}(Q^{k+1})\leq(1-\lambda)^{2}\varsigma+\lambda^{2}\xi.$ (41)
Proof.
Note that
$\mathbb{E}\left[\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})\right]-\mathcal{E}_{\mathcal{D}}(Q^{k+1})$
=$\stackrel{{\scriptstyle(a)}}{{=}}\mathbb{E}\left[\underset{(s,a,s^{\prime})\sim\mathcal{D}}{\mathbb{E}}\,\left[(Q_{\lambda}^{k+1}(s,a)-\hat{\mathcal{B}}Q^{k}(s,a))^{2}\right]\right]-\underset{(s,a,s^{\prime})\sim\mathcal{D}}{\mathbb{E}}\,\left[(Q^{k+1}(s,a)-\hat{\mathcal{B}}Q^{k}(s,a))^{2}\right]$ (42)
=$\stackrel{{\scriptstyle(b)}}{{=}}\underset{(s,a,s^{\prime})\sim\mathcal{D}}{\mathbb{E}}\left[\mathbb{E}\left[(Q_{\lambda}^{k+1}(s,a)-\hat{\mathcal{B}}Q^{k}(s,a))^{2}\right]-(Q^{k+1}(s,a)-\hat{\mathcal{B}}Q^{k}(s,a))^{2}\right]$ (43)
=$\stackrel{{\scriptstyle(c)}}{{=}}\underset{(s,a,s^{\prime})\sim\mathcal{D}}{\mathbb{E}}\left[\mathbb{E}\left[\left(Q_{\lambda}^{k+1}(s,a)\right)^{2}\right]-2\mathbb{E}\left[Q_{\lambda}^{k+1}(s,a)\right]\hat{\mathcal{B}}Q^{k}(s,a)-\left(Q^{k+1}(s,a)\right)^{2}+2Q^{k+1}(s,a)\hat{\mathcal{B}}Q^{k}(s,a)\right]$ (44)
=$\stackrel{{\scriptstyle(d)}}{{=}}\underset{(s,a,s^{\prime})\sim\mathcal{D}}{\mathbb{E}}\left[\mathbb{E}\left[\left(Q_{\lambda}^{k+1}(s,a)\right)^{2}\right]-\left(Q^{k+1}(s,a)\right)^{2}+2\hat{\mathcal{B}}Q^{k}(s,a)\left(Q^{k+1}(s,a)-\mathbb{E}\left[Q_{\lambda}^{k+1}(s,a)\right]\right)\right],$ (45)
where $(a)$ follows from the definition in (4) and (5), $(b)$ follows from exchanging the two expectation, $(c)$ follows by expanding the square, $(d)$ re-arranges the terms. By Proposition 1, we obtain
$\mathbb{E}\left[Q_{\lambda}^{k+1}(s,a)\right]$ =$\stackrel{{\scriptstyle(a)}}{{=}}\underset{s^{\prime}\sim\mathcal{P}_{\mathcal{D}}(s^{\prime}|s,a)}{\mathbb{E}}\left[\frac{1-\lambda}{N(s,a)}\sum_{j=1}^{N(s,a)}\hat{\mathcal{B}}_{s_{j}^{\prime}}Q^{k}(s,a)+\lambda\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)\right]$ (46)
=$\stackrel{{\scriptstyle(b)}}{{=}}(1-\lambda)\underset{s^{\prime}\sim\mathcal{P}_{\mathcal{D}}(s^{\prime}|s,a)}{\mathbb{E}}\left[\frac{1}{N(s,a)}\sum_{j=1}^{N(s,a)}\hat{\mathcal{B}}_{s_{j}^{\prime}}Q^{k}(s,a)\right]+\lambda\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)$ (47)
=$\stackrel{{\scriptstyle(c)}}{{=}}(1-\lambda)\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)+\lambda\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a).$ (48)
where $(a)$ follows from (23), $(b)$ follows from algebraic manipulations, $(c)$ follows by the definition of $\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)$.
Note that $Q^{k+1}(s,a)=\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)$ (by (22) in Proposition 1), and substituting (48) into (45) yields
$\mathbb{E}\left[\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})\right]-\mathcal{E}_{\mathcal{D}}(Q^{k+1})$
$=\underset{(s,a,s^{\prime})\sim\mathcal{D}}{\mathbb{E}}\left[\mathbb{E}\left[\left(Q_{\lambda}^{k+1}(s,a)\right)^{2}\right]-\left(\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)\right)^{2}+2\hat{\mathcal{B}}Q^{k}(s,a)\left(\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)-\left((1-\lambda)\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)+\lambda\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)\right)\right)\right].$ (49)
We further simplified the previous equation as follows
$\mathbb{E}\left[\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})\right]-\mathcal{E}_{\mathcal{D}}(Q^{k+1})$
$=\underset{(s,a,s^{\prime})\sim\mathcal{D}}{\mathbb{E}}\left[\mathbb{E}\left[\left(Q_{\lambda}^{k+1}(s,a)\right)^{2}\right]-\left(\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)\right)^{2}+2\lambda\hat{\mathcal{B}}Q^{k}(s,a)\left(\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)-\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)\right)\right]$ (50)
$=\underset{(s,a,s^{\prime})\sim\mathcal{D}}{\mathbb{E}}\left[\mathbb{E}\left[Q_{\lambda}^{k+1}(s,a)\right]^{2}+\sigma^{2}\left(Q_{\lambda}^{k+1}(s,a)\right)-\left(\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)\right)^{2}+2\lambda\hat{\mathcal{B}}Q^{k}(s,a)\left(\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)-\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)\right)\right],$ (51)
where the last equation follows from the definition of the variance. Substituting (48) into the previous equation yields
$\mathbb{E}\left[\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})\right]-\mathcal{E}_{\mathcal{D}}(Q^{k+1})$
$=\underset{(s,a,s^{\prime})\sim\mathcal{D}}{\mathbb{E}}[\sigma^{2}\left(Q_{\lambda}^{k+1}(s,a)\right)+\left((1-\lambda)\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)+\lambda\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)\right)^{2}$
$\quad\quad\quad\quad\quad-\left(\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)\right)^{2}+2\lambda\hat{\mathcal{B}}Q^{k}(s,a)\left(\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)-\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)\right)]$ (52)
$=\underset{(s,a,s^{\prime})\sim\mathcal{D}}{\mathbb{E}}[\sigma^{2}\left(Q_{\lambda}^{k+1}(s,a)\right)+(1-\lambda)^{2}(\mathcal{B}_{\mathcal{D}}Q^{k}(s,a))^{2}+\lambda^{2}(\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a))^{2}+2\lambda(1-\lambda)\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)$
$\quad\quad\quad\quad\quad-\left(\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)\right)^{2}+2\lambda\hat{\mathcal{B}}Q^{k}(s,a)\left(\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)-\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)\right)]$ (53)
$=\underset{(s,a,s^{\prime})\sim\mathcal{D}}{\mathbb{E}}[\sigma^{2}\left(Q_{\lambda}^{k+1}(s,a)\right)+(\lambda^{2}-2\lambda+1)(\mathcal{B}_{\mathcal{D}}Q^{k}(s,a))^{2}+\lambda^{2}(\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a))^{2}+2\lambda(1-\lambda)\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)$
$\quad\quad\quad\quad\quad-\left(\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)\right)^{2}+2\lambda\hat{\mathcal{B}}Q^{k}(s,a)\left(\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)-\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)\right)],$ (54)
where the previous equations follow by algebraic manipulations. Note that ^$\hat{\mathcal{B}}Q^{k}(s,a)$ is the only term in the previous expression that depends on $s^{\prime}$ (from the domain $\mathcal{D}$ under the expectation, not the samples in $Q_{\lambda}^{k+1}(s,a)$). To proceed, we change the expectation to be w.r.t. $(s,a)\sim\mathcal{D}$
$\mathbb{E}\left[\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})\right]-\mathcal{E}_{\mathcal{D}}(Q^{k+1})$
$=\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[\sigma^{2}\left(Q_{\lambda}^{k+1}(s,a)\right)+(\lambda^{2}-2\lambda+1)(\mathcal{B}_{\mathcal{D}}Q^{k}(s,a))^{2}+\lambda^{2}(\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a))^{2}+2\lambda(1-\lambda)\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)\right.$
$\quad\quad\quad\quad\quad-\left.\left(\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)\right)^{2}+2\lambda\mathbb{E}_{s^{\prime}\sim\mathcal{P}_{\mathcal{D}}(s^{\prime}|s,a)}\left[\hat{\mathcal{B}}Q^{k}(s,a)\right]\left(\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)-\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)\right)\right]$ (55)
$=\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[\sigma^{2}\left(Q_{\lambda}^{k+1}(s,a)\right)+(\lambda^{2}-2\lambda+1)(\mathcal{B}_{\mathcal{D}}Q^{k}(s,a))^{2}+\lambda^{2}(\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a))^{2}+2\lambda(1-\lambda)\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)\right.$
$\quad\quad\quad\quad\quad-\left.\left(\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)\right)^{2}+2\lambda\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)\left(\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)-\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)\right)\right]$ (56)
$=\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[\sigma^{2}\left(Q_{\lambda}^{k+1}(s,a)\right)+\lambda^{2}(\mathcal{B}_{\mathcal{D}}Q^{k}(s,a))^{2}+\lambda^{2}(\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a))^{2}-2\lambda^{2}\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)\right]$ (57)
$=\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[\sigma^{2}\left(Q_{\lambda}^{k+1}(s,a)\right)+\lambda^{2}\left(\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)-\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)\right)^{2}\right],$ (58)
where the previous equations follow by algebraic manipulations. We note that
$\sigma^{2}\left(Q_{\lambda}^{k+1}(s,a)\right)$ =$\stackrel{{\scriptstyle(a)}}{{=}}\sigma^{2}\left(\frac{1-\lambda}{N(s,a)}\sum_{j=1}^{N(s,a)}\hat{\mathcal{B}}_{s_{j}^{\prime}}Q^{k}(s,a)+\lambda\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)\right)$ (59)
=$\stackrel{{\scriptstyle(b)}}{{=}}\sigma^{2}\left(\frac{1-\lambda}{N(s,a)}\sum_{j=1}^{N(s,a)}\hat{\mathcal{B}}_{s_{j}^{\prime}}Q^{k}(s,a)\right)$ (60)
=$\stackrel{{\scriptstyle(c)}}{{=}}\frac{(1-\lambda)^{2}}{N(s,a)}\sigma^{2}\left(\hat{\mathcal{B}}Q^{k}(s,a)\right),$ (61)
where $(a)$ follows by (23) in Proposition 1, $(b)$ is obtained by the theorem of variance of a shifted random variable, $(c)$ is due to the properties of the variance. Thus, substituting the previous expression into (58) yields
$\mathbb{E}\left[\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})\right]-\mathcal{E}_{\mathcal{D}}(Q^{k+1})$ $=\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[\frac{(1-\lambda)^{2}}{N(s,a)}\sigma^{2}\left(\hat{\mathcal{B}}Q^{k}(s,a)\right)+\lambda^{2}\left(\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)-\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)\right)^{2}\right]$ (62)
$\leq\underset{(s,a)\in\mathcal{S}\times\mathcal{A}}{\max}\left[\frac{(1-\lambda)^{2}}{N(s,a)}\sigma^{2}\left(\hat{\mathcal{B}}Q^{k}(s,a)\right)+\lambda^{2}\left(\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)-\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)\right)^{2}\right]$ (63)
$\leq(1-\lambda)^{2}\underset{(s,a)\in\mathcal{S}\times\mathcal{A}}{\max}\left[\frac{\sigma^{2}\left(\hat{\mathcal{B}}Q^{k}(s,a)\right)}{N(s,a)}\right]+\lambda^{2}\underset{(s,a)\in\mathcal{S}\times\mathcal{A}}{\max}\left[\left(\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)-\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)\right)^{2}\right],$ (64)
where the previous inequalities follow from algebraic manipulations. Recall $\xi$ in (12) and $\varsigma$ in (13), we obtain
$\mathbb{E}\left[\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})\right]-\mathcal{E}_{\mathcal{D}}(Q^{k+1})\leq(1-\lambda)^{2}\varsigma+\lambda^{2}\xi.$ (65)
This completes the proof of Theorem 1. ∎
Proof of Theorem 2
Theorem 2 (Worst-Case Performance Bound). Under the hypotheses of Theorem 1, the following bound holds with probability $1-\delta$
$\mathcal{E}_{\mathcal{D}}\left(Q_{\lambda}^{k+1}\right)-\mathcal{E}_{\mathcal{D}}\left(Q^{k+1}\right)\leq(1-\lambda)^{2}\varsigma+\lambda^{2}\xi+\sqrt{\frac{\log\left(1/\delta\right)}{2}}\frac{|\mathcal{S}||\mathcal{A}|}{\sqrt{N}}\frac{\gamma B(1-\lambda)}{(1-\gamma)}\left(8\frac{B(1-\lambda)}{(1-\gamma)}+4\lambda\sqrt{\xi}\right).$ (66)
In particular, when $\lambda$ is selected as in (15), this bound takes the following form
$\mathcal{E}_{\mathcal{D}}\left(Q_{\lambda^{*}}^{k+1}\right)-\mathcal{E}_{\mathcal{D}}\left(Q^{k+1}\right)\leq\frac{\varsigma\xi}{\varsigma+\xi}+\sqrt{\frac{\log\left(1/\delta\right)}{2}}\frac{|\mathcal{S}||\mathcal{A}|}{\sqrt{N}}\frac{\gamma B}{(1-\gamma)}\frac{\xi}{\varsigma+\xi}\!\!\left(\!\frac{8B}{(1-\gamma)}\frac{\xi}{\varsigma+\xi}\!+\!4\frac{\varsigma\sqrt{\xi}}{\varsigma+\xi}\right).$ (67)
Proof.
Note that
$\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})-\mathcal{E}_{\mathcal{D}}(Q^{k+1})$ $=\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})-\mathbb{E}\left[\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})\right]+\mathbb{E}\left[\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})\right]-\mathcal{E}_{\mathcal{D}}(Q^{k+1})$ (68)
$\leq\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})-\mathbb{E}\left[\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})\right]+(1-\lambda)^{2}\varsigma+\lambda^{2}\xi,$ (69)
where the previous inequality follows by Theorem 1. Thus, we are left to bound $\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})-\mathbb{E}\left[\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})\right]$. To proceed, we rely on the following technical lemma.
Lemma 1 (McDiarmid Inequality).
Let $\tau_{1},\cdots,\tau_{n}$ be independent random variables taking on values in a set $H$ and let $c_{1},\cdots,c_{n}$ be positive real constants. If $\varphi:H^{n}\to\mathbb{R}$ satisfies
$\sup_{\tau_{1},\cdots,\tau_{n},\tau_{i}^{\prime}\in H}|\varphi(\tau_{1},\cdots,\tau_{i},\cdots,\tau_{n})-\varphi(\tau_{1},\cdots,\tau_{i}^{\prime},\cdots,\tau_{n})|\leq c_{i},$ (70)
for $1\leq i\leq n$, then it holds that
$P\left(\varphi(\tau_{1},\cdots,\tau_{n})-\mathbb{E}\left[\varphi(\tau_{1},\cdots,\tau_{n})\right]\geq\epsilon\right)\leq\exp\left(\frac{-2\epsilon^{2}}{\sum_{i=1}^{n}c_{i}^{2}}\right).$ (71)
To obtain the similar generalization bound akin to the above, we aim to compute the bound of $\left|\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})-\mathcal{E}_{\mathcal{D}}(\hat{Q}_{\lambda}^{k+1})\right|$, where $\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})$ and $\mathcal{E}_{\mathcal{D}}(\hat{Q}_{\lambda}^{k+1})$ take the random transitions $(\tau_{1},\cdots,\tau_{i},\cdots,\tau_{N})$ and $(\tau_{1},\cdots,\hat{\tau}_{i},\cdots,\tau_{N})$, respectively.
Note that
$\left|\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})-\mathcal{E}_{\mathcal{D}}(\hat{Q}_{\lambda}^{k+1})\right|$
=$\stackrel{{\scriptstyle(a)}}{{=}}\left|\underset{(s,a,s^{\prime})\sim\mathcal{D}}{\mathbb{E}}\left[\left(Q_{\lambda}^{k+1}(s,a)-\hat{\mathcal{B}}Q^{k}(s,a)\right)^{2}-\left(\hat{Q}_{\lambda}^{k+1}(s,a)-\hat{\mathcal{B}}Q^{k}(s,a)\right)^{2}\right]\right|$ (72)
=$\stackrel{{\scriptstyle(b)}}{{=}}\left|\underset{(s,a,s^{\prime})\sim\mathcal{D}}{\mathbb{E}}\left[(Q_{\lambda}^{k+1}(s,a))^{2}-(\hat{Q}_{\lambda}^{k+1}(s,a))^{2}-2Q_{\lambda}^{k+1}(s,a)\hat{\mathcal{B}}Q^{k}(s,a)+2\hat{Q}_{\lambda}^{k+1}(s,a)\hat{\mathcal{B}}Q^{k}(s,a)\right]\right|$ (73)
=$\stackrel{{\scriptstyle(c)}}{{=}}\left|\underset{(s,a,s^{\prime})\sim\mathcal{D}}{\mathbb{E}}\left[(Q_{\lambda}^{k+1}(s,a))^{2}-(\hat{Q}_{\lambda}^{k+1}(s,a))^{2}+2\hat{\mathcal{B}}Q^{k}(s,a)\left(\hat{Q}_{\lambda}^{k+1}(s,a)-Q_{\lambda}^{k+1}(s,a)\right)\right]\right|,$ (74)
where $(a)$ follows from the definition in (4) and (5), $(b)$ and $(c)$ follow from algebraic manipulations.
Note that ^$\hat{\mathcal{B}}Q^{k}(s,a)$ is the only term in the previous expression that depends on $s^{\prime}$ (from the domain $\mathcal{D}$ under the expectation, not the samples in $Q_{\lambda}^{k+1}(s,a)$ or $\hat{Q}_{\lambda}^{k+1}(s,a)$). To proceed, we change the expectation to be w.r.t. $(s,a)\sim\mathcal{D}$
$\left|\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})-\mathcal{E}_{\mathcal{D}}(\hat{Q}_{\lambda}^{k+1})\right|$
$=\left|\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[(Q_{\lambda}^{k+1}(s,a))^{2}-(\hat{Q}_{\lambda}^{k+1}(s,a))^{2}+2\underset{s^{\prime}\sim\mathcal{P}_{\mathcal{D}}(s^{\prime}|s,a)}{\mathbb{E}}\left[\hat{\mathcal{B}}Q^{k}(s,a)\right]\left(\hat{Q}_{\lambda}^{k+1}(s,a)-Q_{\lambda}^{k+1}(s,a)\right)\right]\right|$ (76)
$=\left|\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[(Q_{\lambda}^{k+1}(s,a))^{2}-(\hat{Q}_{\lambda}^{k+1}(s,a))^{2}+2\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)\left(\hat{Q}_{\lambda}^{k+1}(s,a)-Q_{\lambda}^{k+1}(s,a)\right)\right]\right|$ (77)
$=\left|\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[\left(Q_{\lambda}^{k+1}(s,a)-\hat{Q}_{\lambda}^{k+1}(s,a)\right)\left(Q_{\lambda}^{k+1}(s,a)+\hat{Q}_{\lambda}^{k+1}(s,a)-2\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)\right)\right]\right|,$ (78)
where the previous equations follow from algebraic manipulations. Recall that
$Q_{\lambda}^{k+1}(s,a)=\frac{1-\lambda}{N(s,a)}\sum_{j=1}^{N(s,a)}\hat{\mathcal{B}}_{s_{j}^{\prime}}Q^{k}(s,a)+\lambda\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a).$ (79)
Then we obtain
$\left|\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})-\mathcal{E}_{\mathcal{D}}(\hat{Q}_{\lambda}^{k+1})\right|$
=$\stackrel{{\scriptstyle(a)}}{{=}}\left|\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[\left(\frac{1-\lambda}{N(s,a)}(\hat{\mathcal{B}}_{s_{i}^{\prime}}Q^{k}(s,a)-\hat{\mathcal{B}}_{\hat{s}_{i}^{\prime}}Q^{k}(s,a))\right)\left(Q_{\lambda}^{k+1}(s,a)+\hat{Q}_{\lambda}^{k+1}(s,a)-2\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)\right)\right]\right|$ (80)
$\stackrel{{\scriptstyle(b)}}{{\leq}}\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[\left|\left(\frac{1-\lambda}{N(s,a)}(\hat{\mathcal{B}}_{s_{i}^{\prime}}Q^{k}(s,a)-\hat{\mathcal{B}}_{\hat{s}_{i}^{\prime}}Q^{k}(s,a))\right)\right|\cdot\left|\left(Q_{\lambda}^{k+1}(s,a)+\hat{Q}_{\lambda}^{k+1}(s,a)-2\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)\right)\right|\right]$ (81)
$\stackrel{{\scriptstyle(c)}}{{\leq}}\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[\left(\frac{1-\lambda}{N(s,a)}\frac{2\gamma B}{1-\gamma}\right)\left|\left(\frac{1-\lambda}{N(s,a)}\left(\sum_{j=1}^{N(s,a)}\hat{\mathcal{B}}_{s_{j}^{\prime}}Q^{k}(s,a)+\!\!\sum_{j=1}^{N(s,a)}\hat{\mathcal{B}}_{\hat{s}_{j}^{\prime}}Q^{k}(s,a)\right)\!\!+2\lambda\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)-2\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)\right)\right|\right]$ (82)
where $(a)$ is obtained by substituting $Q_{\lambda}^{k+1}(s,a)$, $(b)$ follows from Jensen’s inequality, $(c)$ is obtained by substituting $Q_{\lambda}^{k+1}(s,a)$ and that Assumption 1 and (11) combining with (2) or (3) implies
|^$\left|\hat{\mathcal{B}}_{s_{i}^{\prime}}Q^{k}(s,a)-\hat{\mathcal{B}}_{\hat{s}_{i}^{\prime}}Q^{k}(s,a)\right|\leq\gamma(\frac{B}{1-\gamma}+\frac{B}{1-\gamma})=\frac{2\gamma B}{1-\gamma},\,\forall i$ (83)
|^$\left|\hat{\mathcal{B}}_{s_{j}^{\prime}}Q^{k}(s,a)\right|\leq B+\gamma\frac{B}{1-\gamma}=\frac{B}{1-\gamma},\,\forall j\quad\Rightarrow\quad\left|\sum_{j=1}^{N(s,a)}\hat{\mathcal{B}}_{s_{j}^{\prime}}Q^{k}(s,a)\right|\leq\sum_{j=1}^{N(s,a)}\left|\hat{\mathcal{B}}_{s_{j}^{\prime}}Q^{k}(s,a)\right|\leq N(s,a)\frac{B}{1-\gamma}.$ (84)
By using the previous inequalities and using the triangle inequality, we further have
$\left|\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})-\mathcal{E}_{\mathcal{D}}(\hat{Q}_{\lambda}^{k+1})\right|$ $\leq\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[\left(\frac{1-\lambda}{N(s,a)}\frac{2\gamma B}{1-\gamma}\right)\left((1-\lambda)\frac{2B}{1-\gamma}+\left|2\lambda\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)-2\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)\right|\right)\right]$ (85)
$=\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[\left(\frac{1-\lambda}{N(s,a)}\frac{2\gamma B}{1-\gamma}\right)\left((1-\lambda)\frac{2B}{1-\gamma}+\left|2\lambda\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)-2(1-\lambda+\lambda)\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)\right|\right)\right]$ (86)
$\leq\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[\left(\frac{1-\lambda}{N(s,a)}\frac{2\gamma B}{1-\gamma}\right)\!\!\left((1-\lambda)\frac{2B}{1-\gamma}\!+\!(1-\lambda)\frac{2B}{1-\gamma}\!+\!2\lambda\left|\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)-\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)\right|\right)\right]$ (87)
$\leq\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[\left(\frac{1-\lambda}{N(s,a)}\frac{2\gamma B}{1-\gamma}\right)\!\!\left((1-\lambda)\frac{4B}{1-\gamma}+2\lambda\left|\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)-\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)\right|\right)\right].$ (88)
By definition of the expectation and the fact that $N(s,a)=N\cdot P(s,a)$ (Assumption 2), we have
$\left|\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})-\mathcal{E}_{\mathcal{D}}(\hat{Q}_{\lambda}^{k+1})\right|$
$=\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}P(s,a)\left[\left(\frac{1-\lambda}{N(s,a)}\frac{2\gamma B}{1-\gamma}\right)\left((1-\lambda)\frac{4B}{1-\gamma}+2\lambda\left|\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)-\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)\right|\right)\right]$ (89)
$=\frac{1}{N}\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}\left[\left((1-\lambda)\frac{2\gamma B}{1-\gamma}\right)\left((1-\lambda)\frac{4B}{1-\gamma}+2\lambda\left|\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)-\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)\right|\right)\right]$ (90)
$\leq\frac{1}{N}|\mathcal{S}||\mathcal{A}|\left((1-\lambda)^{2}\frac{8\gamma B^{2}}{(1-\gamma)^{2}}+\lambda(1-\lambda)\frac{4\gamma B}{1-\gamma}\,\,\underset{(s,a)\in\mathcal{S}\times\mathcal{A}}{\max}\left|\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)-\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)\right|\right)$ (91)
$=\frac{|\mathcal{S}||\mathcal{A}|}{N}\left((1-\lambda)^{2}\frac{8\gamma B^{2}}{(1-\gamma)^{2}}+\lambda(1-\lambda)\frac{4\gamma B}{1-\gamma}\sqrt{\xi}\right)$ (92)
$=c$ (93)
where $c$ is the bound as shown in Lemma 1. Lemma 1 implies that
$P\left(\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})-\mathbb{E}\left[\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})\right]\geq\epsilon\right)\leq\exp\frac{-2\epsilon^{2}}{\sum_{i=1}^{N}c^{2}}$ (94)
i.e.,
$P\left(\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})-\mathbb{E}\left[\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})\right]<\epsilon\right)\geq 1-\exp\frac{-2\epsilon^{2}}{Nc^{2}}.$ (95)
Let the right hand side of the previous expression to be $1-\delta$. Then,
$\epsilon$ $=\sqrt{\frac{1}{2}\log(\frac{1}{\delta})N}c$ (96)
$=\sqrt{\frac{1}{2}\log(\frac{1}{\delta})}\frac{|\mathcal{S}||\mathcal{A}|}{\sqrt{N}}\left((1-\lambda)^{2}\frac{8\gamma B^{2}}{(1-\gamma)^{2}}+\lambda(1-\lambda)\frac{4\gamma B}{1-\gamma}\sqrt{\xi}\right).$ (97)
Therefore, the following bound holds with probability $1-\delta$
$\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})-\mathcal{E}_{\mathcal{D}}(Q^{k+1})$ $\leq(1-\lambda)^{2}\varsigma+\lambda^{2}\xi+\sqrt{\frac{1}{2}\log(\frac{1}{\delta})}\frac{|\mathcal{S}||\mathcal{A}|}{\sqrt{N}}\left((1-\lambda)^{2}\frac{8\gamma B^{2}}{(1-\gamma)^{2}}+\lambda(1-\lambda)\frac{4\gamma B}{1-\gamma}\sqrt{\xi}\right)$ (98)
$=(1-\lambda)^{2}\varsigma+\lambda^{2}\xi+\sqrt{\frac{1}{2}\log(\frac{1}{\delta})}\frac{|\mathcal{S}||\mathcal{A}|}{\sqrt{N}}\frac{\gamma B(1-\lambda)}{1-\gamma}\left(8\frac{B(1-\lambda)}{1-\gamma}+4\lambda\sqrt{\xi}\right),$ (99)
which completes the proof of (16).
Then, we further substitute $\lambda^{*}$ defined by (15) into the previous bound and obtain
$\mathcal{E}_{\mathcal{D}}\left(Q_{\lambda^{*}}^{k+1}\right)-\mathcal{E}_{\mathcal{D}}\left(Q^{k+1}\right)\leq\frac{\varsigma\xi}{\varsigma+\xi}+\sqrt{\frac{\log\left(1/\delta\right)}{2}}\frac{|\mathcal{S}||\mathcal{A}|}{\sqrt{N}}\frac{\gamma B}{(1-\gamma)}\frac{\xi}{\varsigma+\xi}\!\!\left(\!\frac{8B}{(1-\gamma)}\frac{\xi}{\varsigma+\xi}\!+\!4\frac{\varsigma\sqrt{\xi}}{\varsigma+\xi}\right).$
This completes the proof of Theorem 2. ∎
Proof of Theorem 3
In this appendix, we provide a corrected version of Theorem 3 together with the comprehensive proof. To proceed, we rely on the following technical lemma.
Lemma 2.
For all iterations $(k=0,1,2,\cdots)$, recall $\xi(Q^{k})$ with respect to $Q^{k}$ as in (12), and let us define
$\xi_{\text{max}}=\sup_{k\in\mathbb{N}}\,\xi(Q^{k}),\quad\sigma_{\text{max}}=\sup_{k\in\mathbb{N}}\,\underset{(s,a)\in\mathcal{S}\times\mathcal{A}}{\max}\mathbb{E}\left[||\frac{1}{N(s,a)}\sum_{j=1}^{N(s,a)}\hat{\mathcal{B}}_{s_{j}^{\prime}}Q_{\lambda}^{k}(s,a)-\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a)||\right].$ (100)
Then, it holds that
$\mathbb{E}\left[\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[||Q_{\lambda}^{k+1}(s,a)-\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a)||_{\infty}\right]\right]\leq(1-\lambda)\sigma_{\text{max}}+\lambda\sqrt{\xi_{\text{max}}},\leavevmode\nobreak\ \forall k\in\mathbb{N}.$ (101)
Proof.
For all iterations $k=0,1,2,\cdots$,
[…]
$\stackrel{{\scriptstyle(a)}}{{\leq}}\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[\mathbb{E}\left[||Q_{\lambda}^{k+1}(s,a)-\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a)||\right]\right]$ (102)
=$\stackrel{{\scriptstyle(b)}}{{=}}\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[\mathbb{E}\left[||\frac{1-\lambda}{N(s,a)}\sum_{j=1}^{N(s,a)}\hat{\mathcal{B}}_{s_{j}^{\prime}}Q_{\lambda}^{k}(s,a)+\lambda\mathcal{B}_{\mathcal{D^{\prime}}}Q_{\lambda}^{k}(s,a)-\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a)||\right]\right]$ (103)
=$\stackrel{{\scriptstyle(c)}}{{=}}\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[\mathbb{E}\left[||(1-\lambda)\left(\frac{1}{N(s,a)}\sum_{j=1}^{N(s,a)}\hat{\mathcal{B}}_{s_{j}^{\prime}}Q_{\lambda}^{k}(s,a)-\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a)\right)+\lambda\left(\mathcal{B}_{\mathcal{D^{\prime}}}Q_{\lambda}^{k}(s,a)-\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a)\right)||\right]\right],$ (104)
where $(a)$ follows from the fact that $||\cdot||_{\infty}\leq||\cdot||$, $(b)$ follows by substituting $Q_{\lambda}^{k+1}(s,a)$ as in (23), $(c)$ follows from algebraic manipulations.
By using the triangle inequality the previous expression yields
[…]
$\leq\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[\mathbb{E}\left[(1-\lambda)||\left(\frac{1}{N(s,a)}\sum_{j=1}^{N(s,a)}\hat{\mathcal{B}}_{s_{j}^{\prime}}Q_{\lambda}^{k}(s,a)-\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a)\right)||+\lambda||\mathcal{B}_{\mathcal{D^{\prime}}}Q_{\lambda}^{k}(s,a)-\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a)||\right]\right]$ (105)
$=\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[(1-\lambda)\mathbb{E}\left[||\left(\frac{1}{N(s,a)}\sum_{j=1}^{N(s,a)}\hat{\mathcal{B}}_{s_{j}^{\prime}}Q_{\lambda}^{k}(s,a)-\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a)\right)||\right]+\lambda||\mathcal{B}_{\mathcal{D^{\prime}}}Q_{\lambda}^{k}(s,a)-\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a)||\right]$ (106)
$=(1-\lambda)\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[\mathbb{E}\left[||\!\!\left(\frac{1}{N(s,a)}\sum_{j=1}^{N(s,a)}\hat{\mathcal{B}}_{s_{j}^{\prime}}Q_{\lambda}^{k}(s,a)-\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a)\right)\!\!||\right]\right]+\lambda\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[||\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a)-\mathcal{B}_{\mathcal{D^{\prime}}}Q_{\lambda}^{k}(s,a)||\right],$ (107)
where the last two equations follow from algebraic manipulations. By definition of $\sigma_{\text{max}}$ and $\xi_{\text{max}}$ we obtain
$\mathbb{E}\left[\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[||Q_{\lambda}^{k+1}(s,a)-\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a)||_{\infty}\right]\right]\leq(1-\lambda)\sigma_{\text{max}}+\lambda\sqrt{\xi_{\text{max}}},\leavevmode\nobreak\ \forall k\in\mathbb{N}.$ (108)
This completes the proof of Lemma 2. ∎
Having introduced Lemma 2, we are in the stage of proving Theorem 3.
Theorem 3 (Convergence). Under the hypotheses of Theorem 1, it holds that
[…]
Proof.
From the triangle inequality we obtain
[…]
Given that $Q^{*}$ is the fixed point of the Bellman optimality operator (2), which is a contraction, we have
[…]
$\leq\mathbb{E}\left[\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[||Q_{\lambda}^{k+1}(s,a)-\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a)||_{\infty}+||\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a)-\mathcal{B}_{\mathcal{D}}Q^{*}(s,a)||_{\infty}\right]\right]$ (110)
$\leq\mathbb{E}\left[\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[||Q_{\lambda}^{k+1}(s,a)-\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a)||_{\infty}+\gamma||Q_{\lambda}^{k}(s,a)-Q^{*}(s,a)||_{\infty}\right]\right]$ (111)
$=\mathbb{E}\left[\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[||Q_{\lambda}^{k+1}(s,a)-\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a)||_{\infty}\right]\right]+\gamma\mathbb{E}\left[\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[||Q_{\lambda}^{k}(s,a)-Q^{*}(s,a)||_{\infty}\right]\right].$ (112)
Unrolling the previous inequality until $Q_{\lambda}^{0}$ yields
[…]
$\leq\mathbb{E}\left[\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[||Q_{\lambda}^{k+1}(s,a)-\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a)||_{\infty}\right]\right]+\gamma\mathbb{E}\left[\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[||Q_{\lambda}^{k}(s,a)-\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k-1}(s,a)||_{\infty}\right]\right]+\cdots$
$+\gamma^{k}\mathbb{E}\left[\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[||Q_{\lambda}^{1}(s,a)-\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{0}(s,a)||_{\infty}\right]\right]+\gamma^{k+1}\mathbb{E}\left[\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[||Q_{\lambda}^{0}(s,a)-Q^{*}(s,a)||_{\infty}\right]\right].$ (113)
Since $Q^{0}$ is the initial Q-function, we have
$Q^{0}(s,a)=Q_{\lambda}^{0}(s,a),\,\forall(s,a)\in\mathcal{S}\times\mathcal{A}.$ (114)
By using Lemma 2 further implies that
[…]
$\leq\left((1-\lambda)\sigma_{\text{max}}+\lambda\sqrt{\xi_{\text{max}}}\right)+\gamma\left((1-\lambda)\sigma_{\text{max}}+\lambda\sqrt{\xi_{\text{max}}}\right)+\cdots+\gamma^{k}\left((1-\lambda)\sigma_{\text{max}}+\lambda\sqrt{\xi_{\text{max}}}\right)$
$\quad+\gamma^{k+1}\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[||Q^{0}(s,a)-Q^{*}(s,a)||_{\infty}\right]$ (115)
$=\gamma^{k+1}\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[||Q^{0}(s,a)-Q^{*}(s,a)||_{\infty}\right]+\frac{1-\gamma^{k+1}}{1-\gamma}\left((1-\lambda)\sigma_{\text{max}}+\lambda\sqrt{\xi_{\text{max}}}\right).$ (116)
This completes the proof of Theorem 3. ∎
Environments
We select five Procgen games (Cobbe et al. 2020) to substantiate our theoretical contributions in this work, whose details are provided below.
Description of Caveflyer (Cobbe et al. 2020).
“The player needs to traverse a complex network of caves to reach the exit. Player movement is reminiscent of the classic Atari game “Asteroids” where the ship can rotate and propel forward or backward along its current axis. The primary reward is granted upon successfully reaching the end of the level, though additional reward can be earned by destroying target objects with the ship’s lasers along the way. The level is fraught with both stationary and moving lethal obstacles, demanding precise navigation and quick reflexes.”
Figure 5: The screenshot of Caveflyer (Cobbe et al. 2020).
Description of Climber (Cobbe et al. 2020).
“The player needs to climb a series of platforms, collecting stars scattered along the path. A small reward is granted for each star collected, with a substantial reward provided for gathering all stars within a level. If every star is collected, the episode terminates. The level is also populated with lethal flying monsters, adding extra challenges to the player’s journey.”
Figure 6: The screenshot of Climber (Cobbe et al. 2020).
Description of Dodgeball (Cobbe et al. 2020).
“Inspired by the Atari game “Berzerk”, the player spawns in a room with a randomly generated configuration of walls and enemies. Contact with a wall results in an immediate game over, terminating the episode. The player moves slowly, allowing for careful navigation throughout the room. Enemies, moving slowly too, throw balls at the player. The player can retaliate by throwing balls as well, but only in the direction they are facing. Once all enemies are eliminated, the player can advance to the unlocked platform, earning a substantial level completion bonus.”
Figure 7: The screenshot of Dodgeball (Cobbe et al. 2020).
Description of Maze (Cobbe et al. 2020).
“The player, embodying a mouse, needs to navigate a maze to locate the sole piece of cheese and obtain a reward. The mazes, generated using Kruskal’s algorithm, vary in size from $3\times 3$ to $25\times 25$, with dimensions uniformly sampled across this range. To navigate the maze, the player can move up, down, left, or right.”
Figure 8: The screenshot of Maze (Cobbe et al. 2020).
Description of Miner (Cobbe et al. 2020).
“Inspired by the game “BoulderDash”, the player (robot) can dig through dirt to navigate the world. The game world is governed by gravity, where dirt supports both boulders and diamonds. Boulders and diamonds fall through free spaces and roll off each other. If either a boulder or a diamond falls on the player, the game terminates immediately. The objective is to collect all the diamonds in the level and then reach the exit. The player earns a small reward for each diamond collected and a huge reward for the successful completion in the level.”
Figure 9: The screenshot of Miner (Cobbe et al. 2020).
Experimental Details
Implementation.
Our implementations as well as the datasets that have been used in this work are based on (Mediratta et al. 2023). Instead of training on a single dataset, this work trains a CQL agent (Kumar et al. 2020) on two different datasets from the source and target domains. These two domains contain distinct environment levels, e.g., $[0,99]$ for the source domain and $[100,199]$ for the target domain.
Hyperparameters.
We adhere to the hyperparameters set by the original implementation in (Mediratta et al. 2023), with the sole exception of assigning distinct weights to each of the two dataset. The main hyperparameters employed in this work are summarized in Table 5.
Table 5: Experimental hyperparameters.
Hyperparameters Value
Target domain levels $[100,199]$
Source domain levels $[0,99]$, $[25,124]$, $[50,149]$
Number of target samples ($N$) $1000,2500,4000$
Number of source samples $40000$
Weight ($\lambda$) $\{0,0.2,0.4,0.5,0.6,0.8,1\}$
Number of episodes for evaluation $500$
Learning rate $0.0005$
Batch size $256$
Neural network hidden size $256$
Discount factor ($\gamma$) $0.99$
CQL conservativeness constant ($\alpha$) $4$
Gradient norm clip $0.1$
Computing Infrastructure
All numerical experiments were implemented on a workstation with an NVIDIA GeForce RTX 3070 GPU, Intel Core i9-10900 CPU at 2.80 GHz.
Code
The codes will be made available upon the publication of this work.