### v3

# Provable Domain Adaptation for Offline Reinforcement Learning with Limited Samples

Abstract
Offline reinforcement learning (RL) learns effective policies from a static target dataset. The performance of state-of-the-art offline RL algorithms notwithstanding, it relies on the quality and size of the target dataset and it degrades if limited samples in the target dataset are available, which is often the case in real-world applications. To address this issue, domain adaptation that leverages auxiliary samples from related source datasets (such as simulators) can be beneficial. However, establishing the optimal way to trade off the source and target datasets while ensuring provably theoretical guarantees remains an open challenge. To the best of our knowledge, this paper proposes the first framework that theoretically explores the impact of the weights assigned to each dataset on the performance of offline RL. In particular, we establish performance bounds and the existence of an optimal weight, which can be computed in closed form under simplifying assumptions. We also provide algorithmic guarantees in terms of convergence to a neighborhood of the optimum. Notably, these results depend on the quality of the source dataset and the number of samples from the target dataset. Our empirical results on the well-known Procgen benchmark substantiate our theoretical contributions.
1 Introduction
Deep reinforcement learning (RL) has demonstrated impressive performance in a wide variety of applications, such as strategy games (Mnih et al., 2013; 2015), robotics (Levine et al., 2016; Duan et al., 2016), and recommender systems (Afsar et al., 2022; Lin et al., 2023). RL aims to learn an optimal policy that maximizes the expected cumulative reward. To achieve this goal, the RL agent learns and improves the policy by actively interacting with the environment. However, this poses a critical challenge for the real-world applications of RL, as interactions with the real-world can be significantly dangerous and expensive (Kumar et al., 2020; Levine et al., 2020; Chen et al., 2024). In this context, offline RL merges as a promising alternative framework for the real-world applications of RL where the agent learns effective policies from a static and previously-collected dataset.
Recent advances in offline RL algorithms have shown outstanding performance across a diverse array of problems and datasets (Fujimoto et al., 2019a; Kumar et al., 2020; Kostrikov et al., 2021; Chen et al., 2021). Nevertheless, their effectiveness depends on the quality and size of the dataset. More concretely, it is worth noting that even state-of-the-art (SOTA) offline RL algorithms like BCQ (Fujimoto et al., 2019a), CQL (Kumar et al., 2020), IQL (Kostrikov et al., 2021) , DT (Chen et al., 2021) demonstrate poor performance given a small offline RL dataset, as training on a small number of samples may lead to overfitting (Fu et al., 2019; Kumar et al., 2019). In this work, we are interested in offline RL that learns from a static dataset with limited samples.
1.1 Related Work
Offline reinforcement learning with dataset distillation.
Dataset distillation (Wang et al., 2018) proposes a framework for synthesizing a smaller and more efficient dataset by minimizing the gradient discrepancy of the samples from the original dataset and the distilled dataset. Synthetic (Light et al., 2024) is the first work that applies dataset distillation to offline RL and achieves comparable performance with the original large offline RL dataset. Specifically, it synthesizes a small distilled dataset by minimizing the gradient matching loss between the original offline RL dataset and the synthetic dataset. However, generating the synthetic dataset necessitates access to the original large offline RL dataset, which is often impractical in real-world scenarios. In particular, this work focuses on scenarios where only a limited number of samples are accessible.
Offline reinforcement learning with domain adaptation.
Figure 1: Schematic: the target dataset has limited samples (red), whereas the source dataset has unlimited samples (green) with dynamics gap from the target domain. How to strike a proper balance between the two datasets?
To avoid overfitting by learning from the limited target dataset, domain adaptation techniques (Redko et al., 2020; Farahani et al., 2021) proposes to leverage an auxiliary large source dataset (such as simulators) with unlimited samples. H2O (Niu et al., 2022) assumes access to an unrestricted simulator and it requires knowing the dynamics of both datasets, and the process of training on simulators still requires interacting with the environment online. On the other hand, ORIS (Hou et al., 2024) proposes to generate a new (source) dataset from the simulators, where a generative adversarial network (GAN) model is employed to approximate the state distribution of the original target dataset. Starting from the initial state provided by GAN, the new (source) dataset is generated by interacting with the simulator and reweighted by an additional discriminator model.
Given the unlimited source dataset and the limited target dataset, however, striking the proper balance between the two datasets remains a challenging problem in offline RL (see Figure 1). Natural solutions involve either combining both datasets equally or using only one of them. For instance, H2O consistently treats both datasets with equal weights. On the other hand, ORIS suggests always assigning a lower weight to the source dataset than to the target dataset, as the dynamics gap could negatively impact the performance of the offline RL. It is worth pointing out that the weights assigned in these cases lack guarantees. Accordingly, this work focuses on exploring the impact of the weights assigned to the two datasets on the performance of offline RL, particularly from the provably theoretical perspectives.
Offline reinforcement learning with dynamics gap.
Dynamics gap in the domain adaptation for offline RL is widely acknowledged as a significant challenge. DARC (Eysenbach et al., 2020) and DARA (Liu et al., 2022) train a dynamics gap-related penalty by minimizing the divergence between the real and simulator trajectory distributions, and combine it with the simulator reward during the online/offline training of simulators. However, to our knowledge, no prior work has theoretically investigated the impact of the dynamics gap on selecting the weights assigned to the limited target dataset and the unlimited source dataset.
1.2 Main Contributions
The main contributions of this paper are summarized as follows.
  • To the best of our knowledge, this is the first work that proposes an algorithm-agnostic framework of domain adaptation for offline RL, which exhibits a provably theoretical trade-off between the number of samples in a limited target dataset and the discrepancy (or dynamics gap) between the target and source domains.
  • We establish the (expected and worst-case) performance bounds and the convergence to a neighborhood of the optimum within our framework. We further identify under simplifying assumptions the existence of an optimal weight for balancing the two datasets, which is typically not one of the trivial choices: treating both datasets equally or using either dataset only. All theoretical guarantees and the optimal weight will depend on the quality of the source dataset (refer to the dynamics gap) and the size of the target dataset (i.e., number of samples).
  • A series of numerical experiments conducted on the offline Procgen Benchmark (Mediratta et al., 2023) substantiate our theoretical contributions.
2 Balancing Target and Source Datasets
In this section, we consider the mathematical formalism of offline RL (Levine et al., 2020), namely Markov Decision Process (MDP) (Sutton & Barto, 2018). MDP is defined by a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},P,r,\rho,\gamma,T)$, where $\mathcal{S}$ and $\mathcal{A}$ are finite state and action spaces, $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to[0,1]$ denotes the transition probability that describes the dynamics of the system, $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ denotes the reward function that evaluates the quality of the action, $\rho:\mathcal{S}\to[0,1]$ denotes the initial state distribution, $\gamma$ represents the discount factor, and $T\in\{0,1,2,\cdots\}$ defines the horizon length.
A policy is represented by the probability distribution over actions conditioned on states. In each state $s$, the agent selects an action $a$ based on the policy $\pi(a|s)$, and generates the next state $s^{\prime}$. The tuple $(s,a,s^{\prime})$ is referred to as the transition data. Offline RL considers employing a behavior policy $\pi_{\beta}$ to collect an offline and static dataset with $N$ transitions ($N\in\mathbb{N}_{+}$), i.e., $\hat{\mathcal{D}}=\{(s_{i},a_{i},s_{i}^{\prime})\}_{i=1}^{N}$. The transitions in the dataset $\hat{\mathcal{D}}$ are collected from a domain $\mathcal{D}$. The goal of RL is to learn an optimal policy $\pi^{*}$ that maximizes the expected discounted cumulative reward, i.e.,
$\pi^{*}=\operatorname*{arg\,max}_{\pi}\,\underset{\begin{subarray}{c}s_{0}\sim\rho,\,a_{t}\sim\pi(a_{t}|s_{t})\\ s_{t+1}\sim P(s_{t+1}|s_{t},a_{t})\end{subarray}}{\mathbb{E}}\left[\sum_{t=0}^{T}\gamma^{t}r(s_{t},a_{t})\right].$ (1)
In the context of offline RL, the policy needs to be learned from the dataset $\hat{\mathcal{D}}$ exclusively. Offline RL algorithms (Fujimoto et al., 2019a; Levine et al., 2020; Kostrikov et al., 2021) typically train an action-value function (or Q-function) by minimizing the temporal difference (TD) error iteratively. To be formal, let $\mathcal{B}$ be a Bellman operator. This operator takes different forms depending on the specific algorithm considered. For instance, in Q-learning type methods the operator takes the form
$\mathcal{B}Q^{k}(s,a)=r(s,a)+\gamma\underset{s^{\prime}\sim P(s^{\prime}|s,a)}{\mathbb{E}}\left[\max_{a^{\prime}}Q^{k}(s^{\prime},a^{\prime})\right],$ (2)
whereas for actor-critic methods it takes the form
$\mathcal{B}^{\pi}Q^{k}(s,a)=r(s,a)+\gamma\underset{\begin{subarray}{c}s^{\prime}\sim P(s^{\prime}|s,a),\,a^{\prime}\sim\pi(a^{\prime}|s^{\prime})\end{subarray}}{\mathbb{E}}\left[Q^{k}(s^{\prime},a^{\prime})\right].$ (3)
Given any transition $(s,a,s^{\prime})$, let us define the TD error for any $Q$-function at the step $k$ as
$\mathcal{E}(Q,(s,a,s^{\prime})):=\left(Q(s,a)-\hat{\mathcal{B}}_{s^{\prime}}Q^{k}(s,a)\right)^{2},$ (4)
where ^ssubscript^$\hat{\mathcal{B}}_{s^{\prime}}$ denotes the stochastic approximation of the Bellman operator, namely a version of (2) or (3) where the expectation is replaced by evaluating the random variable at a single of realization $s^{\prime}$. Note that we have dropped the dependence on $k$ on the left hand side of the above expression for simplicity. We further define the expected TD error in the domain $\mathcal{D}$ as
$\mathcal{E}_{\mathcal{D}}(Q)=\underset{(s,a,s^{\prime})\sim\mathcal{D}}{\mathbb{E}}\,\left[\mathcal{E}(Q,(s,a,s^{\prime}))\right].$ (5)
With these definitions, one can define the iterations of a large class of offline RL algorithms through the following optimization problem
$Q^{k+1}=\operatorname*{arg\,min}_{Q}\,\mathcal{E}_{\mathcal{D}}(Q),\,\forall k\in\mathbb{N}.$ (6)
It is worth pointing out that the specific forms of (2) and (3) can result in poor performance in offline RL attributed to the issues with bootstrapping from out-of-distribution (OOD) actions (Fujimoto et al., 2019b; Kumar et al., 2019; 2020; Levine et al., 2020). This typically leads to an overestimation of the Q-value. To avoid this overestimation, prior works consider solely using in-distribution state-action pairs to maintain the Q-function (Fujimoto et al., 2019a), or constraining the learned policy to remain closely aligned with the behavior policy (Levine et al., 2020). These algorithms, however, can also be written as (5) by introducing modifications to the Bellman operators.
Note that the expectation in (5) poses a challenge in solving problem (6): it requires visiting every transition infinite times. In practice, one defines the empirical version of the TD error $\mathcal{E}_{\mathcal{D}}(Q)$ in (5),
$\mathcal{E}_{\hat{\mathcal{D}}}(Q)=\frac{1}{N}\sum_{i=1}^{N}\mathcal{E}(Q,(s_{i},a_{i},s_{i}^{\prime})),$ (7)
where the samples are from the dataset $\hat{\mathcal{D}}$. Then, the offline RL algorithm is defined as the minimization of the stochastic approximation of problem (6)
$\hat{Q}^{k+1}=\operatorname*{arg\,min}_{Q}\,\mathcal{E}_{\hat{\mathcal{D}}}(Q),\,\forall k\in\mathbb{N}.$ (8)
It has been widely demonstrated that SOTA offline RL algorithms, such as BCQ (Fujimoto et al., 2019a), CQL (Kumar et al., 2020), IQL (Kostrikov et al., 2021), and DT (Chen et al., 2021), are capable of solving problem (6) given sufficient transition samples from the domain $\mathcal{D}$. Nevertheless, this assumption may not be realizable in practice, e.g., healthcare (Tang et al., 2022) and autonomous driving (Pan et al., 2017), where data collection is challenging. In this regime, $\hat{Q}^{k+1}$ generally demonstrates poor performance in approximating $Q^{k+1}$, as training on a small number of samples can lead to overfitting (Fu et al., 2019; Kumar et al., 2019).
On the other hand, in certain applications, one can rely on simulators (or related datasets) that provide a larger number of samples $\mathcal{D}^{\prime}=\{(s_{j},a_{j},s_{j}^{\prime})\}_{j=1}^{N^{\prime}}$ with $N^{\prime}\gg N$. It is worth noting that, in general, $\mathcal{D}^{\prime}$ will differ from $\mathcal{D}$ in terms of the state distribution and transition probabilities. Similar to (5), we define
$\mathcal{E}_{\mathcal{D}^{\prime}}(Q)=\underset{(s_{j},a_{j},s_{j}^{\prime})\sim\mathcal{D}^{\prime}}{\mathbb{E}}\,\left[\mathcal{E}(Q,(s_{j},a_{j},s_{j}^{\prime}))\right].$ (9)
To be precise, one still needs to work with the empirical TD error on the dataset $\mathcal{D}^{\prime}$ since estimating (9) is infeasible. Note however, that since $N^{\prime}\gg N$ the generalization error due to having a finite number of samples in the source dataset $\mathcal{D}^{\prime}$ is negligible. For this reason and to simplify the exposition we choose to assume that one can compute (9), and denote by $\mathcal{D}^{\prime}$ both the source dataset and the source domain with a slight abuse of notation.
Prior works such as H2O (Niu et al., 2022) and ORIS (Hou et al., 2024) propose to improve the performance of the limited target dataset, by combining with a far larger source dataset. Significantly, H2O consistently treats both datasets equally, whereas ORIS always assigns a lower weight to the source dataset. In this paper, we explore a more general approach to combine both datasets
$Q_{\lambda}^{k+1}=\operatorname*{arg\,min}_{Q}\,(1-\lambda)\mathcal{E}_{\hat{\mathcal{D}}}(Q)\!+\!\lambda\mathcal{E}_{\mathcal{D}^{\prime}}(Q),\,\forall k\in\mathbb{N},$ (10)
where $\lambda\in[0,1]$ denotes the weight that trades off $\mathcal{E}_{\hat{\mathcal{D}}}(Q)$ from the limited target dataset and $\mathcal{E}_{\mathcal{D}^{\prime}}(Q)$ from the unlimited source dataset. Particularly, (10) recovers H2O when $\lambda=0.5$ and yields ORIS when $\lambda<0.5$. Indeed, $\lambda\approx 1$ prioritizes the minimization of the TD error corresponding to the source dataset. This approach is suitable in cases where the source dataset is similar to the target (or coming from the same domain in an extreme case). On the other hand, $\lambda\approx 0$ focuses on minimizing the TD error in the target dataset. This method is appropriate for scenarios where data is abundant (very large $N$) or where the discrepancies between the target and source datasets are too large.
Given these observations, it is expected that different values of $\lambda$ can attain the optimal performance, depending on the interplay between the available number of target samples and the discrepancies between the datasets. The following section formalizes this expectation.
3 Performance and Convergence Guarantees
We start this section by discussing the necessary assumptions to develop our theoretical results concerning the generalization of the solution to problem (10).
Assumption 1.
There exists $B>0$ such that, for any $(s,a)\in\mathcal{S}\times\mathcal{A}$, $|r(s,a)|<B$.
Assumption 1 is common in the literature (Azar et al., 2017; Wei et al., 2020; Zhang et al., 2021). In particular, in the case of finite state-action spaces, it is always possible to design the reward to avoid the possibility of being unbounded. Further notice that this assumption along with the fact that a geometric series with ratio $\gamma\in[0,1)$ converges to $1/(1-\gamma)$ implies that
$\max_{(s,a)\in\mathcal{S}\times\mathcal{A}}Q(s,a)\leq\frac{B}{1-\gamma}.$ (11)
To proceed, we denote by $P_{\mathcal{D^{\prime}}}(s,a)$ and $P_{\mathcal{D}}(s,a)$ the probability of a state-action pair $(s,a)$ from the source and target domains. Let $N(s,a)\in\left\{1,2,\ldots,N\right\}$ be the number of the specific $(s,a)$ pairs in the target dataset $\hat{\mathcal{D}}$ and define $P_{\hat{\mathcal{D}}}(s,a)=N(s,a)/N$.
Assumption 2.
For any $(s,a)\in\mathcal{S}\times\mathcal{A}$, $P_{\mathcal{D}}(s,a)$, $P_{\mathcal{D}^{\prime}}(s,a)$ and $P_{\hat{\mathcal{D}}}(s,a)$ are positive, and there exist constants $\beta_{u}\geq\beta_{l}>0$ such that the following bounded ratios holds:
$\frac{P_{\hat{\mathcal{D}}}(s,a)}{P_{\mathcal{D}^{\prime}}(s,a)}\in[\beta_{l},\beta_{u}],\quad\frac{P_{\hat{\mathcal{D}}}(s,a)}{P_{\mathcal{D}}(s,a)}\in[\beta_{l},\beta_{u}],\quad\frac{P_{\mathcal{D}^{\prime}}(s,a)}{P_{\mathcal{D}}(s,a)}\in[\beta_{l},\beta_{u}].$
It is worth noting that the transition sample $\tau$ are independent and identically distributed (i.i.d.), as the dataset $\hat{\mathcal{D}}$ is shuffled and the transition data are sampled i.i.d.. Alternatively, one can think of the dataset as random samples from the occupancy measure, followed by a random transition.
Building upon Assumption 2, we derive the analytical expressions for $Q^{k+1}(s,a)$ and $Q_{\lambda}^{k+1}(s,a)$ for any $(s,a)\in\mathcal{S}\times\mathcal{A}$, which we formally present in the following proposition.
Proposition 1.
Let Assumption 2 hold. Recall the empirical Bellman operator $\hat{\mathcal{B}}$ in (4). Denote by $\mathcal{B}_{\mathcal{D}}$ ($\mathcal{B}_{\mathcal{D^{\prime}}}$) the Bellman operator in (2) or (3) in which $s^{\prime}$ follows the transition probability of the domain $\mathcal{D}$ ($\mathcal{D^{\prime}}$). Recall that $Q^{k+1}$ and $Q_{\lambda}^{k+1}$ represent the solutions to (6) and (10). Denote by $N$ the number of samples in the target dataset $\hat{\mathcal{D}}$ and $N(s,a)\in\{1,2,\cdots,N\}$ the amount of $(s,a)$ pairs in $\hat{\mathcal{D}}$. At each iteration $(k=1,2,3,\cdots)$, it holds that
$Q^{k+1}(s,a)=\mathcal{B}_{\mathcal{D}}Q^{k}(s,a),\,\forall(s,a)\in\mathcal{S}\times\mathcal{A},$ (12)
$Q_{\lambda}^{k+1}(s,a)=\frac{\frac{1-\lambda}{N}\sum_{j=1}^{N(s,a)}\hat{\mathcal{B}}_{s_{j}^{\prime}}Q^{k}(s,a)+\lambda P_{\mathcal{D}^{\prime}}(s,a)\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)}{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)+\lambda P_{\mathcal{D}^{\prime}}(s,a)},\,\forall(s,a)\in\mathcal{S}\times\mathcal{A}.$ (13)
We proceed by exploring the performance of $Q_{\lambda}^{k+1}(s,a)$, the solution to our problem of interest (10). To do so, we next define two quantities that will play important roles in the theoretical guarantees of $Q_{\lambda}^{k+1}(s,a)$. Start by defining the discrepancy or dynamic gap between the two domains
$\xi=\underset{(s,a)\in\mathcal{S}\times\mathcal{A}}{\max}\left[\left(\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)-\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)\right)^{2}\right].$ (14)
where, for notation simplicity, we omit the dependence of $\xi$ on the iteration $k$. Indeed, if the transition probabilities in the source and target domains match, the quantity $\xi$ in (14) is zero. We also define a measure of the variability of the target dataset
$\varsigma=\underset{\begin{subarray}{c}(s,a)\in\mathcal{S}\times\mathcal{A}\end{subarray}}{\max}\left[\underset{\hat{\mathcal{D}}}{\max}\Bigg{[}\frac{\sigma_{\hat{s}^{\prime}}^{2}\left(\hat{\mathcal{B}}_{\hat{s}^{\prime}}Q^{k}(s,a)\right)}{N(s,a)}\Bigg{]}\right],$ (15)
where $\sigma_{\hat{s}^{\prime}}^{2}\left(\hat{\mathcal{B}}_{\hat{s}^{\prime}}Q^{k}(s,a)\right)$ denotes the variance of the empirical Bellman operator. The term $\varsigma$ above is the maximum normalized variance across all state-action pairs in the spaces. It is significant to note that the number of samples to keep the value $\varsigma$ constant is proportional to the variance.
With these assumptions, definitions and the analytical expression for $Q_{\lambda}^{k+1}$, we are in conditions of establishing the bound of the expected value of TD error for $Q_{\lambda}^{k+1}$ over the target domain.
Theorem 1 (Expected Performance Bound).
Let Assumptions 1 and 2 hold. Recall $Q^{k+1}$ in (12), $Q_{\lambda}^{k+1}$ in (13), $\xi$ in (14), and $\varsigma$ in (15). For all iterations $(k=1,2,3,\cdots)$, it holds that
$\underset{\begin{subarray}{c}N(s,a),\,\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\left[\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})\right]-\mathcal{E}_{\mathcal{D}}(Q^{k+1})\leq\left(\frac{1-\lambda}{1-\lambda+\lambda/\beta_{u}}\right)^{2}\varsigma+\left(\frac{\lambda}{(1-\lambda)\beta_{l}+\lambda}\right)^{2}\xi.$ (16)
The previous theorem not only establishes the expected performance bound of $Q_{\lambda}^{k+1}$ but also implies the bias-variance trade-off sought by combining the two datasets with different weights in (10). Indeed, the optimal weight $\lambda^{*}$ that minimizes the right hand side of the above bound is discussed formally by the following corollaries.
Corollary 1.
With Theorem 1 holding, the optimal weight $\lambda^{*}$ that minimizes the bound in (16) is 00 when $\varsigma=0$ and $1$ when $\xi=0$.
Corollary 2.
With Theorem 1 holding and $\beta_{l}=\beta_{u}=\beta>0$, the optimal weight $\lambda^{*}$ that minimizes the bound in (16) takes the form
$\lambda^{*}=\frac{\beta^{2}\varsigma}{\beta^{2}\varsigma+\beta\xi}.$ (17)
It is crucial to highlight that both corollaries above recover the intuition that the target dataset with no variation (or the number of samples in the target dataset is very large), i.e., $\varsigma\approx 0$, encourages to consider the target dataset only in (10), i.e., $\lambda^{*}=0$. On the other hand, when the two domains are close ($\xi\approx 0$), the optimal value of $\lambda$ is one, suggesting that one should use the source dataset solely. Although intuitive, the expected performance bound in Theorem 1 is insufficient to claim any generalization guarantees as the tails of the distribution could be heavy. We address this concern in the next theorem by providing the generalization bound (worst-case performance bound).
Theorem 2 (Worst-Case Performance Bound).
Under the hypotheses of Theorem 1, the following bound holds with probability at least $1-\delta$
$\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})-\mathcal{E}_{\mathcal{D}}(Q^{k+1})$
$\leq\left(\frac{1-\lambda}{1-\lambda+\lambda/\beta_{u}}\right)^{2}\varsigma+\left(\frac{\lambda}{(1-\lambda)\beta_{l}+\lambda}\right)^{2}\xi$
$\quad\,+\sqrt{\frac{1}{2}\log(\frac{1}{\delta})}\frac{|\mathcal{S}||\mathcal{A}|}{\sqrt{N}}\left(\frac{8\gamma B^{2}}{(1-\gamma)^{2}}\frac{\beta_{u}(1-\lambda)^{2}}{(1-\lambda)\beta_{l}^{2}+\lambda\beta_{l}}+\frac{4\gamma B}{1-\gamma}\frac{\lambda(1-\lambda)\sqrt{\xi}}{(1-\lambda)\beta_{l}^{2}+\lambda\beta_{l}}\right).$ (18)
The above theorem provides the worst-case bound of solving (10), which demonstrates the bias-variance trade-off by the two datasets with different weights as well. Most importantly, both the expected and worst-case performance bounds, as shown in (16) and (2), imply that the optimal trade-off between the source and target datasets is not always trivial, meaning that $\lambda^{*}$ may not belong to $\{0,0.5,1\}$. The optimal trade-off for the expected performance depends on the number of samples in the target dataset (corresponding to $\varsigma$), the discrepancy (or dynamics gap) between the two domains (corresponding to $\xi$), and the bounds of $P_{\hat{\mathcal{D}}}(s,a)/P_{\mathcal{D}^{\prime}}(s,a)$. In addition to $\varsigma$, $\xi$, $\beta_{l}$ and $\beta_{u}$, the optimal weight for the worst-case performance will depend on more factors such as the reward bound $B$, the size of the state and action spaces $|\mathcal{S}|$ and $|\mathcal{A}|$.
Having established the performance bounds of solving (10), we are in the stage of providing the convergence guarantee. We formalize it in the next theorem, which relies on the following two quantities: the maximum of the dynamics gap over all iterations and the maximum of the empirical standard deviation over all iterations.
$\xi_{\text{max}}=\sup_{k\in\mathbb{N}}\,\xi(Q^{k}),\,\sigma_{\text{max}}=\sup_{k\in\mathbb{N}}\underset{\begin{subarray}{c}(s,a)\in\mathcal{S}\times\mathcal{A}\end{subarray}}{\max}\left[\!\underset{\hat{\mathcal{D}}}{\max}\!\left[||\!\left(\frac{\sum_{j=1}^{N(s,a)}\hat{\mathcal{B}}_{s_{j}^{\prime}}Q_{\lambda}^{k}(s,a)}{N(s,a)}-\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a)\right)\!||\right]\right].$ (19)
Theorem 3 (Convergence).
Under the hypotheses of Theorem 1, it holds that
$\underset{\begin{subarray}{c}N(s,a),\,\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\left[\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[||Q_{\lambda}^{k+1}(s,a)-Q^{*}(s,a)||_{\infty}\right]\right]\leq$
$\gamma^{k+1}\!\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\!\left[||Q^{0}(s,a)-Q^{*}(s,a)||_{\infty}\right]\!+\!\frac{1-\gamma^{k+1}}{1-\gamma}\!\!\left(\frac{1-\lambda}{1-\lambda+\lambda/\beta_{u}}\sigma_{\text{max}}\!+\!\frac{\lambda}{(1-\lambda)\beta_{l}+\lambda}\sqrt{\xi_{\text{max}}}\right).$ (20)
The previous theorem implies that the solution $Q_{\lambda}^{k+1}$ of solving (10) is guaranteed in expectation to converge to a neighborhood of the optimal $Q$-function, i.e., $Q^{*}$ as $k\to\infty$. This neighborhood is presented as follows
$\mathcal{C}=\frac{1}{1-\gamma}\left(\frac{1-\lambda}{1-\lambda+\lambda/\beta_{u}}\sigma_{\text{max}}+\frac{\lambda}{(1-\lambda)\beta_{l}+\lambda}\sqrt{\xi_{\text{max}}}\right).$ (21)
In addition to the discount factor $\gamma$, the neighborhood $\mathcal{C}$ depends on the weight $\lambda$, the maximal dynamics gap $\xi_{\text{max}}$, and the maximal empirical standard deviation $\sigma_{\text{max}}$, and the bounds $\beta_{l}$ and $\beta_{u}$.
4 Numerical Experiments
Although this work primarily focuses on theoretical analyses, we present in this section a series of numerical experiments that demonstrate the performance of solving (10) under different weight $\lambda$ and validate the corresponding theoretical contributions in the previous section.
4.1 Environmental Setup
We consider the well-known offline Procgen Benchmark (Mediratta et al., 2023) that is often used to assess the domain adaptation/generalization of offline RL. We select five games from Procgen: Caveflyer, Climber, Dodgeball, Maze, Miner. The environmental details including the introduction of the five games are deferred to Appendix B.
Recall that our framework is algorithm-agnostic, meaning that various SOTA offline RL algorithms such as BCQ, CQL, and IQL can apply. In this work, we select CQL as a representative algorithm, as it typically demonstrates a promising and robust performance across a variety of offline RL tasks (Kumar et al., 2020; Fu et al., 2020). To ensure a fair comparison, we retain all hyperparameters consistent, e.g., batch size, learning rate, and network size, and solely change the weight assigned to each dataset. The implementation details are provided in Appendix C.
(a) Caveflyer
(b) Climber
(c) Dodgeball
(d) Maze
(e) Miner
Figure 2: The performance of offline RL with domain adaptation across five Procgen games. The source dataset contains $40000$ samples from levels $[0,99]$, while the target dataset comprises $1000$ samples from levels $[100,199]$. We consider seven weights, $\lambda\in\{0,0.2,0.4,0.5,0.6,0.8,1.0\}$, to trade off the source and target datasets with the star marking the optimal weight.
(a) Caveflyer
(b) Climber
(c) Dodgeball
(d) Maze
(e) Miner
Figure 3: The performance of offline RL with domain adaptation across five Procgen games. The target dataset comprises $1000$ samples from levels $[100,199]$, and three source datasets are considered, each containing $40000$ samples from levels $[0,99]$ (green, $\xi_{1}$), $[25,124]$ (blue, $\xi_{2}$), and $[50,149]$ (red, $\xi_{3}$), respectively. Seven weights, $\lambda\in\{0,0.2,0.4,0.5,0.6,0.8,1.0\}$, are evaluated to trade off the source and target datasets with the star marking the optimal weight for each $\xi$.
(a) Caveflyer
(b) Climber
(c) Dodgeball
(d) Maze
(e) Miner
Figure 4: The performance of offline RL with domain adaptation across five Procgen games. The source dataset comprises $40000$ samples from levels $[50,149]$, and target datasets from levels $[100,199]$ are considered with three different sample sizes: $N=1000$ (green), $N=2500$ (blue), and $N=4000$ (red). Seven weights, $\lambda\in\{0,0.2,0.4,0.5,0.6,0.8,1.0\}$, are evaluated to trade off the source and target datasets with the star marking the optimal weight for each $N$.
4.2 Results
Recall that the worst-case performance bound in (2) can be considerably conservative, particularly when the size of the state and action spaces are very large. Therefore, this subsection focuses on the expected performance bound (16) as well as its corresponding corollaries. It is crucial to note that the variance $\varsigma$ and the dynamics gap $\xi$ in (16) are challenging to measure or estimate precisely, as it requires access to the entire source and target domains. This is not possible in our problem of interest. Nevertheless, one can still investigate how these factors influence the expected bound (16), which provides insights into the expected performance of offline RL. Specifically, we examine the impact of each of $\varsigma$, $\xi$, and $\lambda$ on the expected performance. To achieve this, we vary one of these factors at a time while keeping the other two parameters constant.
Note that Procgen employs procedural content generation to create adaptive levels upon episode reset. Each level corresponds to a specific seed (non-negative integer) and has distinct dynamics and visual appearances. We consider the source and target domains to be with levels $[0,99]$ and $[100,199]$, respectively. Recall that the target dataset is expected to contain significantly fewer samples than the source dataset, i.e., $N\ll N^{\prime}$. Typically, $N^{\prime}$ is considered to be ten times larger than $N$. Therefore, we set $N=1000$ and sample from levels $[100,199]$ for the target dataset, and set $N^{\prime}=40000$ and sample from levels $[0,99]$ for the source dataset. We denote by $\xi_{1}$ the discrepancy (or dynamics gap) between levels $[0,99]$ and $[100,199]$.
Impact of $\lambda$.
We consider seven $\lambda$ from $\{0,0.2,0.4,0.5,0.6,0.8,1\}$, where $\lambda\in\{0,0.5,1\}$ represents the three trivial choices. However, (16) and Corollary 2 reveal that the optimal weight that minimizes the expected performance bound may not be the three trivial choices. Indeed, our experiments on the Procgen, as depicted in Figure 2, validates this claim by the observation that only one out of five games has the optimal weight to be the trivial choice. In addition, Miner shows that all three trivial choices of $\lambda$ achieve worse performance compared to other weights.
Impact of $\xi$.
We consider three source datasets of the same size but from different domains: levels $[0,99]$, levels $[25,124]$, $[50,149]$. Similar to $\xi_{1}$, let $\xi_{2}$ and $\xi_{3}$ represent the discrepancies (or dynamics gap) between levels $[25,124]$ and $[100,199]$, and between levels $[50,149]$ and $[100,199]$, respectively. Thus, we obtain $\xi_{1}\geq\xi_{2}\geq\xi_{3}$, as more overlap between the levels of the two datasets demonstrates smaller discrepancies between them. The bound (16) in Theorem 1 implies that the expected performance improves as $\xi$ decreases. Our numerical results of five games across three different $\xi$ values are summarized in Table 1 (see Appendix C.1), which supports the implication from the bound (16).
Impact of $\varsigma$.
As the normalized variance $\varsigma$ decreases with $N(s,a)$ (see (15)) and given positive proportional relationship between $N$ and $N(s,a)$, a larger $N$ practically leads to a smaller $\varsigma$ (in expectation). The bound (16) in Theorem 1 implies that the expected performance improves as $\varsigma$ decreases. Our numerical results of five games across three different values of $N$ are summarized in Table 2 (see Appendix C.1), which validates the implication from the bound (16).
Recall that Corollary 2 explicitly demonstrates that a smaller $\xi$ drives $\lambda^{*}$ closer to 1, while a smaller $\varsigma$ shifts $\lambda^{*}$ closer to 0. We substantiate this through empirical evidence. $i)$ To explore how $\lambda^{*}$ varies with $\xi$, we fix $N=1000$ in each of the five Procgen games and select three different source datasets with $\xi_{1}\geq\xi_{2}\geq\xi_{3}$. We then implement this using seven values of $\lambda\in\{0,0.2,0.4,0.5,0.6,0.8,1\}$. Our numerical results are summarized in Figure 3 and Table 3 (see Appendix C.1). Both results demonstrate that the optimal weight $\lambda^{*}$ within $\{0,0.2,0.4,0.5,0.6,0.8,1\}$ increases (closer to 1) as $\xi$ decreases. This substantiates Corollary 2 and the intuition that greater emphasis should be placed on the source dataset when its discrepancy from the target domain is smaller. In the extreme case, where the source and target domains are identical, one should consider the source dataset only. $ii)$ To explore how $\lambda^{*}$ varies with $\varsigma$, we fix $\xi_{3}$ in each of the five Procgen games. Since a larger $N$ corresponds to a smaller $\varsigma$, we select three different values of $N\in\{1000,2500,4000\}$. It is worth noting that $N^{\prime}$ is consistently maintained at ten times larger than $N$, ensuring that the source dataset always contains significantly more samples than the target dataset. We then implement this using seven values of $\lambda\in\{0,0.2,0.4,0.5,0.6,0.8,1\}$. Our numerical results are summarized in Figure 4 and Table 4 (see Appendix C.1). Both results demonstrate that the optimal weight $\lambda^{*}$ within $\{0,0.2,0.4,0.5,0.6,0.8,1\}$ decreases (closer to 0) as $N$ increases (i.e., $\varsigma$ decreases). This validates Corollary 2 and the intuition that greater emphasis should be placed on the target dataset when it contains more samples. In the extreme case where the target dataset contains an infinite number of samples, it becomes optimal to rely exclusively on the target dataset.
5 Conclusion
Offline RL is highly dependent on the size of the target dataset. Even state-of-the-art offline RL algorithms often lack performance guarantees under a limited number of samples. To tackle offline RL with limited samples, domain adaptation can be employed which considers related source datasets e.g., simulators that typically offer unlimited (or a very large number of) samples. To the best of our knowledge, we propose in this paper the first framework that theoretically and experimentally explores the domain adaptation for offline RL with limited samples. Specifically, we establish the expected and worst-case performance bounds, as well as a convergence neighborhood under our framework. Moreover, this work provides the optimal weight for trading off the unlimited source dataset and the limited target dataset. It demonstrates that the optimal weight is not necessarily one of the trivial choices: using either dataset only or combining the two datasets equally. We conduct a series of numerical experiments on the renowned benchmark Procgen. Our numerical results validate the theoretical contributions in this work.
Supplementary Materials
The following content was not necessarily subject to peer review.
In the supplementary materials, we provide the omitted proofs of our theoretical contributions together with experimental details in this work.
Appendix A Omitted Proofs
A.1 Proof of Proposition 1
Proposition 1. Let Assumption 2 hold. Recall the empirical Bellman operator $\hat{\mathcal{B}}$ in (4). Denote by $\mathcal{B}_{\mathcal{D}}$ ($\mathcal{B}_{\mathcal{D^{\prime}}}$) the Bellman operator in (2) or (3) in which $s^{\prime}$ follows the transition probability of the domain $\mathcal{D}$ ($\mathcal{D^{\prime}}$). Recall that $Q^{k+1}$ and $Q_{\lambda}^{k+1}$ represent the solutions to (6) and (10). Denote by $N$ the number of samples in the target dataset $\hat{\mathcal{D}}$ and $N(s,a)\in\{1,2,\cdots,N\}$ the amount of $(s,a)$ pairs in $\hat{\mathcal{D}}$. At each iteration $(k=1,2,3,\cdots)$, it holds that
$Q^{k+1}(s,a)=\mathcal{B}_{\mathcal{D}}Q^{k}(s,a),\,\forall(s,a)\in\mathcal{S}\times\mathcal{A},$ (22)
$Q_{\lambda}^{k+1}(s,a)=\frac{\frac{1-\lambda}{N}\sum_{j=1}^{N(s,a)}\hat{\mathcal{B}}_{s_{j}^{\prime}}Q^{k}(s,a)+\lambda P_{\mathcal{D}^{\prime}}(s,a)\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)}{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)+\lambda P_{\mathcal{D}^{\prime}}(s,a)},\,\forall(s,a)\in\mathcal{S}\times\mathcal{A}.$ (23)
Proof.
Note that (12) is well-known in RL, however, we provide its proof here for completeness. For any $Q$, we note that
$\mathcal{E}_{\mathcal{D}}(Q)$ =$\stackrel{{\scriptstyle(a)}}{{=}}\underset{(s,a,s^{\prime})\sim\mathcal{D}}{\mathbb{E}}\left[\left(Q(s,a)-\hat{\mathcal{B}}_{s^{\prime}}Q^{k}(s,a)\right)^{2}\right]$ (24)
=$\stackrel{{\scriptstyle(b)}}{{=}}\sum_{s,a}P_{D}(s,a)\sum_{s^{\prime}\sim\mathcal{D}}P_{\mathcal{D}}(s^{\prime}\mid s,a)\left(Q(s,a)-\hat{\mathcal{B}}_{s^{\prime}}Q^{k}(s,a)\right)^{2},$ (25)
where $(a)$ holds by definitions (4) and (5), $(b)$ follows from the definition of conditional expectation.
The derivative of $\mathcal{E}_{\mathcal{D}}(Q)$ w.r.t. $Q(s,a),\forall(s,a)\in\mathcal{S}\times\mathcal{A}$ is given by
$\frac{\partial\mathcal{E}_{\mathcal{D}}(Q)}{\partial Q(s,a)}$ =$\stackrel{{\scriptstyle(a)}}{{=}}2P_{\mathcal{D}}(s,a)\sum_{s^{\prime}}P_{\mathcal{D}}(s^{\prime}|s,a)\left(Q(s,a)-\hat{\mathcal{B}}_{s^{\prime}}Q^{k}(s,a)\right)$ (26)
=$\stackrel{{\scriptstyle(b)}}{{=}}2P_{\mathcal{D}}(s,a)\left(Q(s,a)-\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)\right),$ (27)
where $(a)$ follows by taking the derivative, leveraging the fact that each $Q(s,a)$ is only present in exactly one term of the summation (27), $(b)$ follows from the facts that $Q(s,a)$ is independent of $s^{\prime}$ and by definition $\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)=\mathbb{E}_{s^{\prime}\sim P_{\mathcal{D}}(s^{\prime}|s,a)}\left[\hat{\mathcal{B}}_{s^{\prime}}Q^{k}(s,a)\right]$ (see (2) or (3)). Since the objective (6) is strongly convex, its minimizer $Q^{k+1}$ is the unique point that satisfies
$\frac{\partial\mathcal{E}_{\mathcal{D}}(Q^{k+1})}{\partial Q(s,a)}=0,\,\forall(s,a)\in\mathcal{S}\times\mathcal{A}.$ (28)
Combining the previous equation with (27) and the fact that $P_{\mathcal{D}}(s,a)>0$ (by Assumption 2) yields
$Q^{k+1}(s,a)=\mathcal{B}_{\mathcal{D}}Q^{k}(s,a),\,\forall(s,a)\in\mathcal{S}\times\mathcal{A}.$ (29)
This completes the proof of (12) in Proposition 1.
We now turn our attention to proving (13). Note that
$(1-\lambda)\mathcal{E}_{\hat{\mathcal{D}}}(Q)+\lambda\mathcal{E}_{\mathcal{D}^{\prime}}(Q)$ =$\stackrel{{\scriptstyle(a)}}{{=}}(1-\lambda)\frac{1}{N}\sum_{i=1}^{N}\left(Q(s_{i},a_{i})-\hat{\mathcal{B}}_{s^{\prime}}Q^{k}(s_{i},a_{i})\right)^{2}+\lambda\mathcal{E}_{\mathcal{D}^{\prime}}(Q)$ (30)
=$\stackrel{{\scriptstyle(b)}}{{=}}(1-\lambda)\frac{1}{N}\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}\sum_{j=1}^{N(s,a)}\left(Q(s,a)-\hat{\mathcal{B}}_{s_{j}^{\prime}}Q^{k}(s,a)\right)^{2}+\lambda\mathcal{E}_{\mathcal{D}^{\prime}}(Q),$ (31)
where $(a)$ follows from the definitions in (4) and (7), $(b)$ re-arranges (30) by summing first the $N(s,a)$ samples of the next state $s^{\prime}$ corresponding to a given $(s,a)$ pair.
The derivative of $(1-\lambda)\mathcal{E}_{\hat{\mathcal{D}}}(Q)+\lambda\mathcal{E}_{\mathcal{D}^{\prime}}(Q)$ w.r.t. $Q(s,a),\forall(s,a)\in\mathcal{S}\times\mathcal{A}$ is given by
$\frac{\partial\left((1-\lambda)\mathcal{E}_{\hat{\mathcal{D}}}(Q)+\lambda\mathcal{E}_{\mathcal{D}^{\prime}}(Q)\right)}{\partial Q(s,a)}$
=$\stackrel{{\scriptstyle(a)}}{{=}}(1-\lambda)\frac{2}{N}\sum_{j=1}^{N(s,a)}\left(Q(s,a)-\hat{\mathcal{B}}_{s_{j}^{\prime}}Q^{k}(s,a)\right)+2\lambda P_{\mathcal{D^{\prime}}}(s,a)\left(Q(s,a)-\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)\right)$ (32)
=$\stackrel{{\scriptstyle(b)}}{{=}}(1-\lambda)\frac{2}{N}\left(N(s,a)Q(s,a)-\sum_{j=1}^{N(s,a)}\hat{\mathcal{B}}_{s_{j}^{\prime}}Q^{k}(s,a)\right)+2\lambda P_{\mathcal{D^{\prime}}}(s,a)\left(Q(s,a)-\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)\right),$ (33)
where the first term in $(a)$ is obtained by taking the derivative,leveraging the fact that each $Q(s,a)$ is only present in exactly one term of the summation (33), and the second term in $(a)$ is analogous to the derivation of $\partial\mathcal{E}_{\mathcal{D}}(Q)/\partial Q(s,a)$ (see (27)) replacing $\mathcal{D}$ by $\mathcal{D^{\prime}}$, (b) is due to the fact that $\sum_{j=1}^{N(s,a)}Q(s,a)=N(s,a)Q(s,a)$.
Similar to (28), since $Q_{\lambda}^{k+1}$ is the unique solution to the convex problem (10), the gradient of the objective zero. Equating the right nad side of (33) to zero follows that
$Q_{\lambda}^{k+1}(s,a)\!\left((1-\lambda)P_{\hat{\mathcal{D}}}(s,a)+\lambda P_{\mathcal{D^{\prime}}}(s,a)\right)\!=\!\frac{1-\lambda}{N}\sum_{j=1}^{N(s,a)}\!\!\hat{\mathcal{B}}_{s_{j}^{\prime}}Q^{k}(s,a)+\lambda P_{\mathcal{D^{\prime}}}(s,a)\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a),$ (34)
where in the above expression we use $P_{\hat{\mathcal{D}}}(s,a)=N(s,a)/N$ as in Assumption 2. We note that $(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)+\lambda P_{\mathcal{D}^{\prime}}(s,a)>0$ as $P_{\hat{\mathcal{D}}}(s,a)>0,P_{\mathcal{D^{\prime}}}(s,a)>0$ (by Assumption 2) and $\lambda\in[0,1]$. Then we further have
$Q_{\lambda}^{k+1}(s,a)=\frac{\frac{1-\lambda}{N}\sum_{j=1}^{N(s,a)}\hat{\mathcal{B}}_{s_{j}^{\prime}}Q^{k}(s,a)+\lambda P_{\mathcal{D}^{\prime}}(s,a)\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)}{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)+\lambda P_{\mathcal{D}^{\prime}}(s,a)},\,\forall(s,a)\in\mathcal{S}\times\mathcal{A}.$ (35)
This completes the proof of Proposition 1. ∎
A.2 Proof of Theorem 1
Theorem 1 (Expected Performance Bound). Let Assumptions 1 and 2 hold. Recall $Q^{k+1}$ in (12), $Q_{\lambda}^{k+1}$ in (13), $\xi$ in (14), and $\varsigma$ in (15). For all iterations $(k=1,2,3,\cdots)$, it holds that
$\underset{\begin{subarray}{c}N(s,a),\,\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\left[\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})\right]-\mathcal{E}_{\mathcal{D}}(Q^{k+1})\leq\left(\frac{1-\lambda}{1-\lambda+\lambda/\beta_{u}}\right)^{2}\varsigma+\left(\frac{\lambda}{(1-\lambda)\beta_{l}+\lambda}\right)^{2}\xi.$ (36)
Proof.
We start by writing $\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})-\mathcal{E}_{D}(Q^{k+1})$ using the definitions in (4) and (5).
$\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})-\mathcal{E}_{D}(Q^{k+1})=\underset{(s,a,s^{\prime})\sim\mathcal{D}}{\mathbb{E}}\Big{[}(Q_{\lambda}^{k+1}(s,a)-\hat{\mathcal{B}}_{s^{\prime}}Q^{k}(s,a))^{2}-(Q^{k+1}(s,a)-\hat{\mathcal{B}}_{s^{\prime}}Q^{k}(s,a))^{2}\Big{]}.$ (37)
Expanding the squares in the above expression follows that
$\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})-\mathcal{E}_{D}(Q^{k+1})$
$=\underset{(s,a,s^{\prime})\sim\mathcal{D}}{\mathbb{E}}\Big{[}(Q_{\lambda}^{k+1}(s,a))^{2}-(Q^{k+1}(s,a))^{2}+2\hat{\mathcal{B}}_{s^{\prime}}Q^{k}(s,a)\left(Q^{k+1}(s,a)-Q^{k+1}_{\lambda}(s,a)\right)\Big{]}.$ (38)
Conditioning on $(s,a)$, using the fact that only $\hat{B}_{s^{\prime}}Q^{k}(s,a)$ depends on $s^{\prime}$, and that by definition $\mathbb{E}_{s^{\prime}\sim P_{\mathcal{D}}(s^{\prime}|s,a)}\left[\hat{\mathcal{B}}_{s^{\prime}}Q^{k}(s,a)\right]=\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)$, the above equation reduces to
$\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})-\mathcal{E}_{D}(Q^{k+1})$
$=\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\Big{[}(Q_{\lambda}^{k+1}(s,a))^{2}-(Q^{k+1}(s,a))^{2}+2\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)\left(Q^{k+1}(s,a)-Q^{k+1}_{\lambda}(s,a)\right)\Big{]}.$ (39)
Replacing $Q^{k+1}(s,a)$ with $\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)$ (by (12) in Proposition 1) in the previous equation yields
$\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})-\mathcal{E}_{D}(Q^{k+1})$
$=\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\Big{[}(Q_{\lambda}^{k+1}(s,a))^{2}-\left(\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)\right)^{2}+2\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)\left(\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)-Q^{k+1}_{\lambda}(s,a)\right)\Big{]}$ (40)
$=\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\Big{[}(Q_{\lambda}^{k+1}(s,a))^{2}+\left(\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)\right)^{2}-2\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)Q^{k+1}_{\lambda}(s,a)\Big{]}.$ (41)
We next take the expectation on (41) with respect to $\hat{s}^{\prime}$ (from the dataset $\hat{\mathcal{D}}$), using the fact that only $Q_{\lambda}^{k+1}(s,a)$ depends on $\hat{s}^{\prime}$ and the conditional expectation it follows thatSince the dataset $\hat{\mathcal{D}}$ is sampled i.i.d., we can swap the two expectations above to obtain
$\underset{\begin{subarray}{c}\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\left[\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})\right]-\mathcal{E}_{D}(Q^{k+1})$
$=\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\Bigg{[}\underset{\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)}{\mathbb{E}}\Big{[}\left(Q_{\lambda}^{k+1}(s,a)\right)^{2}\Big{]}+\left(\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)\right)^{2}$
$\quad\quad\quad\quad\quad-2\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)\Bigg{(}\underset{\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)}{\mathbb{E}}\Big{[}Q_{\lambda}^{k+1}(s,a)\Big{]}\Bigg{)}\Bigg{]}.$ (42)
Using the definition of the variance of a random variable, we rewrite the first term in the right hand side of the above expression
$\underset{\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)}{\mathbb{E}}\Big{[}\left(Q_{\lambda}^{k+1}(s,a)\right)^{2}\Big{]}=\sigma_{\hat{s}^{\prime}}^{2}\left(Q_{\lambda}^{k+1}(s,a)\right)+\left(\underset{\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)}{\mathbb{E}}\Big{[}Q_{\lambda}^{k+1}(s,a)\Big{]}\right)^{2}.$ (43)
Substituting the previous expression in (42) yields
$\underset{\begin{subarray}{c}\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\left[\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})\right]-\mathcal{E}_{D}(Q^{k+1})$
$=\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\Bigg{[}\sigma_{\hat{s}^{\prime}}^{2}\left(Q_{\lambda}^{k+1}(s,a)\right)+\left(\underset{\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)}{\mathbb{E}}\Big{[}Q_{\lambda}^{k+1}(s,a)\Big{]}\right)^{2}+\left(\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)\right)^{2}$
$\quad\quad\quad\quad\quad-2\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)\Bigg{(}\underset{\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)}{\mathbb{E}}\Big{[}Q_{\lambda}^{k+1}(s,a)\Big{]}\Bigg{)}\Bigg{]}.$ (44)
Note that the last three terms on the right hand side of the above expression are the square of a difference. Hence, the previous equation reduces to
$\underset{\begin{subarray}{c}\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\left[\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})\right]-\mathcal{E}_{D}(Q^{k+1})$
$=\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\Bigg{[}\sigma_{\hat{s}^{\prime}}^{2}\left(Q_{\lambda}^{k+1}(s,a)\right)+\left(\underbrace{\underset{\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)}{\mathbb{E}}\Big{[}Q_{\lambda}^{k+1}(s,a)\Big{]}-\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)}_{U_{1}}\right)^{2}\Bigg{]}.$ (45)
We next work on $\sigma_{\hat{s}^{\prime}}^{2}\left(Q_{\lambda}^{k+1}(s,a)\right)$ and $U_{1}$ separately by focusing on $\sigma_{\hat{s}^{\prime}}^{2}\left(Q_{\lambda}^{k+1}(s,a)\right)$ first. For any fixed $N(s,a)$, we note that
$\sigma_{\hat{s}^{\prime}}^{2}\left(Q_{\lambda}^{k+1}(s,a)\right)$ =$\stackrel{{\scriptstyle(a)}}{{=}}\sigma_{\hat{s}^{\prime}}^{2}\left(\frac{\frac{1-\lambda}{N}\sum_{j=1}^{N(s,a)}\hat{\mathcal{B}}_{\hat{s}_{j}^{\prime}}Q^{k}(s,a)+\lambda P_{\mathcal{D}^{\prime}}(s,a)\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)}{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)+\lambda P_{\mathcal{D}^{\prime}}(s,a)}\right)$ (46)
=$\stackrel{{\scriptstyle(b)}}{{=}}\sigma_{\hat{s}^{\prime}}^{2}\left(\frac{\frac{1-\lambda}{N}\sum_{j=1}^{N(s,a)}\hat{\mathcal{B}}_{\hat{s}_{j}^{\prime}}Q^{k}(s,a)}{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)+\lambda P_{\mathcal{D}^{\prime}}(s,a)}\right)$ (47)
$=\sigma_{\hat{s}^{\prime}}^{2}\left(\frac{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)\frac{1}{N(s,a)}\sum_{j=1}^{N(s,a)}\hat{\mathcal{B}}_{\hat{s}_{j}^{\prime}}Q^{k}(s,a)}{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)+\lambda P_{\mathcal{D}^{\prime}}(s,a)}\right)$ (48)
=$\stackrel{{\scriptstyle(c)}}{{=}}\left(\frac{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)}{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)+\lambda P_{\mathcal{D}^{\prime}}(s,a)}\right)^{2}\frac{\sigma_{\hat{s}^{\prime}}^{2}\left(\hat{\mathcal{B}}_{\hat{s}^{\prime}}Q^{k}(s,a)\right)}{N(s,a)},$ (49)
where $(a)$ follows from (13) in Proposition 1, $(b)$ is obtained by the theorem of variance of a shifted random variable, since $\lambda P_{\mathcal{D}^{\prime}}(s,a)\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)$ does not depend on $\hat{s}^{\prime}$, $(c)$ is due to $\sigma^{2}(cX)=c^{2}\sigma^{2}(X)$ where $X$ denotes any random variable and $c$ is a constant.
We now turn our attention to $U_{1}$. For any fixed $N(s,a)$, we obtain
$U_{1}$ =$\stackrel{{\scriptstyle(a)}}{{=}}\underset{\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)}{\mathbb{E}}\left[\frac{\frac{1-\lambda}{N}\sum_{j=1}^{N(s,a)}\hat{\mathcal{B}}_{\hat{s}_{j}^{\prime}}Q^{k}(s,a)+\lambda P_{\mathcal{D}^{\prime}}(s,a)\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)}{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)+\lambda P_{\mathcal{D}^{\prime}}(s,a)}\right]-\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)$ (50)
=$\stackrel{{\scriptstyle(b)}}{{=}}\frac{\frac{1-\lambda}{N}N(s,a)\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)+\lambda P_{\mathcal{D}^{\prime}}(s,a)\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)}{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)+\lambda P_{\mathcal{D}^{\prime}}(s,a)}-\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)$ (51)
=$\stackrel{{\scriptstyle(c)}}{{=}}\frac{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)+\lambda P_{\mathcal{D}^{\prime}}(s,a)\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)}{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)+\lambda P_{\mathcal{D}^{\prime}}(s,a)}-\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)$ (52)
=$\stackrel{{\scriptstyle(d)}}{{=}}\frac{\lambda P_{\mathcal{D}^{\prime}}(s,a)\left(\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)-\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)\right)}{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)+\lambda P_{\mathcal{D}^{\prime}}(s,a)},$ (53)
where $(a)$ follows from (13) in Proposition 1, $(b)$ follows by the definition of $\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)$ (see (2) or (3)), $(c)$ follows by the definition $P_{\hat{\mathcal{D}}}(s,a)=N(s,a)/N$, and $(d)$ follows from combining and canceling terms.
Now, substituting the expressions for $\sigma_{\hat{s}^{\prime}}^{2}\left(Q_{\lambda}^{k+1}(s,a)\right)$ and $U_{1}$ from (49) and (53) into (45) results in
$\underset{\begin{subarray}{c}\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\left[\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})\right]-\mathcal{E}_{\mathcal{D}}(Q^{k+1})$
$=\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\Bigg{[}\underbrace{\left(\frac{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)}{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)+\lambda P_{\mathcal{D}^{\prime}}(s,a)}\right)^{2}}_{U_{2}}\frac{\sigma_{\hat{s}^{\prime}}^{2}\left(\hat{\mathcal{B}}_{\hat{s}^{\prime}}Q^{k}(s,a)\right)}{N(s,a)}+$
$\quad\quad\quad\quad\quad\quad\quad\quad\underbrace{\left(\frac{\lambda P_{\mathcal{D}^{\prime}}(s,a)}{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)+\lambda P_{\mathcal{D}^{\prime}}(s,a)}\right)^{2}}_{U_{3}}\left(\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)-\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)\right)^{2}\Bigg{]}.$ (54)
Dividing both the numerator and denominator by $P_{\hat{\mathcal{D}}}(s,a)$ in $U_{2}$ and by $P_{\mathcal{D}^{\prime}}(s,a)$ in $U_{3}$ yields
$\underset{\begin{subarray}{c}\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\left[\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})\right]-\mathcal{E}_{\mathcal{D}}(Q^{k+1})$
$=\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\Bigg{[}\left(\frac{1-\lambda}{1-\lambda+\lambda P_{\mathcal{D}^{\prime}}(s,a)/P_{\hat{\mathcal{D}}}(s,a)}\right)^{2}\frac{\sigma_{\hat{s}^{\prime}}^{2}\left(\hat{\mathcal{B}}_{\hat{s}^{\prime}}Q^{k}(s,a)\right)}{N(s,a)}+$
$\quad\quad\quad\quad\quad\quad\quad\quad\left(\frac{\lambda}{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)/P_{\mathcal{D}^{\prime}}(s,a)+\lambda}\right)^{2}\left(\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)-\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)\right)^{2}\Bigg{]}.$ (55)
Assumption 2 implies that both $P_{\hat{\mathcal{D}}}(s,a)/P_{\mathcal{D}^{\prime}}(s,a)$ and $P_{\mathcal{D}^{\prime}}(s,a)/P_{\hat{\mathcal{D}}}(s,a)$ are bounded, i.e.,
$\beta_{l}\leq\frac{P_{\hat{\mathcal{D}}}(s,a)}{P_{\mathcal{D}^{\prime}}(s,a)}\leq\beta_{u},\quad\frac{1}{\beta_{u}}\leq\frac{P_{\mathcal{D}^{\prime}}(s,a)}{P_{\hat{\mathcal{D}}}(s,a)}\leq\frac{1}{\beta_{l}}.$ (56)
Substituting the previous expression into (55) yields
$\underset{\begin{subarray}{c}\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\left[\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})\right]-\mathcal{E}_{\mathcal{D}}(Q^{k+1})$
$\leq\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\Bigg{[}\left(\frac{1-\lambda}{1-\lambda+\lambda/\beta_{u}}\right)^{2}\frac{\sigma_{\hat{s}^{\prime}}^{2}\left(\hat{\mathcal{B}}_{\hat{s}^{\prime}}Q^{k}(s,a)\right)}{N(s,a)}$ (57)
$\quad\quad\quad\quad\quad+\left(\frac{\lambda}{(1-\lambda)\beta_{l}+\lambda}\right)^{2}\left(\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)-\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)\right)^{2}\Bigg{]}$ (58)
$=\left(\frac{1-\lambda}{1-\lambda+\lambda/\beta_{u}}\right)^{2}\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\Bigg{[}\frac{\sigma_{\hat{s}^{\prime}}^{2}\left(\hat{\mathcal{B}}_{\hat{s}^{\prime}}Q^{k}(s,a)\right)}{N(s,a)}\Bigg{]}$
$\quad+\left(\frac{\lambda}{(1-\lambda)\beta_{l}+\lambda}\right)^{2}\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\Bigg{[}\left(\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)-\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)\right)^{2}\Bigg{]}.$ (59)
By the definition of $\xi$ as in (14), the above expression is equivalent to
$\underset{\begin{subarray}{c}\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\left[\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})\right]-\mathcal{E}_{\mathcal{D}}(Q^{k+1})$
$\leq\left(\frac{1-\lambda}{1-\lambda+\lambda/\beta_{u}}\right)^{2}\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\Bigg{[}\frac{\sigma_{\hat{s}^{\prime}}^{2}\left(\hat{\mathcal{B}}_{\hat{s}^{\prime}}Q^{k}(s,a)\right)}{N(s,a)}\Bigg{]}+\left(\frac{\lambda}{(1-\lambda)\beta_{l}+\lambda}\right)^{2}\xi.$ (60)
To recover the objective function of Theorem 1 (as in (16)), we take the expectation on the previous expression with respect to $N(s,a)$ (from the dataset $\hat{\mathcal{D}}$) and using the fact that only $Q_{\lambda}^{k+1}$ depends on $N(s,a)$ yields
$\underset{\begin{subarray}{c}N(s,a),\,\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\left[\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})\right]-\mathcal{E}_{\mathcal{D}}(Q^{k+1})$
$\leq\left(\frac{1-\lambda}{1-\lambda+\lambda/\beta_{u}}\right)^{2}\underset{\begin{subarray}{c}(s,a)\sim\mathcal{D}\end{subarray}}{\mathbb{E}}\left[\underset{\begin{subarray}{c}N(s,a)\end{subarray}}{\mathbb{E}}\Bigg{[}\frac{\sigma_{\hat{s}^{\prime}}^{2}\left(\hat{\mathcal{B}}_{\hat{s}^{\prime}}Q^{k}(s,a)\right)}{N(s,a)}\Bigg{]}\right]+\left(\frac{\lambda}{(1-\lambda)\beta_{l}+\lambda}\right)^{2}\xi$ (61)
$\leq\left(\frac{1-\lambda}{1-\lambda+\lambda/\beta_{u}}\right)^{2}\underset{\begin{subarray}{c}(s,a)\in\mathcal{S}\times\mathcal{A}\end{subarray}}{\max}\left[\underset{\hat{\mathcal{D}}}{\max}\Bigg{[}\frac{\sigma_{\hat{s}^{\prime}}^{2}\left(\hat{\mathcal{B}}_{\hat{s}^{\prime}}Q^{k}(s,a)\right)}{N(s,a)}\Bigg{]}\right]+\left(\frac{\lambda}{(1-\lambda)\beta_{l}+\lambda}\right)^{2}\xi.$ (62)
By the definition of $\varsigma$ in (15), we obtain
$\underset{\begin{subarray}{c}N(s,a),\,\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\left[\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})\right]-\mathcal{E}_{\mathcal{D}}(Q^{k+1})\leq\left(\frac{1-\lambda}{1-\lambda+\lambda/\beta_{u}}\right)^{2}\varsigma+\left(\frac{\lambda}{(1-\lambda)\beta_{l}+\lambda}\right)^{2}\xi.$ (63)
This completes the proof of Theorem 1. ∎
A.3 Proof of Corollary 1
Corollary 1. With Theorem 1 holding, the optimal weight $\lambda^{*}$ that minimizes the bound in (16) is 00 when $\varsigma=0$ and $1$ when $\xi=0$.
Proof.
We start by taking the derivative of the right hand side of (16) with respect to $\lambda$, i.e.,
$2\varsigma\left(\frac{1-\lambda}{1-\lambda+\lambda/\beta_{u}}\right)\left(\frac{-\left(1-\lambda+\lambda/\beta_{u}\right)-(1-\lambda)(1/\beta_{u}-1)}{\left(1-\lambda+\lambda/\beta_{u}\right)^{2}}\right)$ (64)
$+2\xi\left(\frac{\lambda}{(1-\lambda)\beta_{l}+\lambda}\right)\frac{(1-\lambda)\beta_{l}+\lambda-\lambda(1-\beta_{l})}{\left((1-\lambda)\beta_{l}+\lambda\right)^{2}}.$ (65)
By combining and canceling the terms the previous expression reduces to
$2\varsigma\frac{(\lambda-1)/\beta_{u}}{\left(1-\lambda+\lambda/\beta_{u}\right)^{3}}+2\xi\frac{\lambda\beta_{l}}{\left((1-\lambda)\beta_{l}+\lambda\right)^{3}}.$ (66)
Notice that the optimal weight $\lambda^{*}$ that minimizes the bound in (16) should be chosen such that the derivative above is zero, i.e.,
$0\displaystyle 2\varsigma\frac{(\lambda^{*}-1)/\beta_{u}}{\left(1-\lambda^{*}+\lambda^{*}/\beta_{u}\right)^{3}}+2\xi\frac{\lambda^{*}\beta_{l}}{\left((1-\lambda^{*})\beta_{l}+\lambda^{*}\right)^{3}}=$ 0 . (67)
where both $\left(1-\lambda^{*}+\lambda^{*}/\beta_{u}\right)$ and $\left((1-\lambda^{*})\beta_{l}+\lambda^{*}\right)$ are positive for any $\lambda^{*}\in[0,1]$ due to that $\beta_{u}\geq\beta_{l}>0$.
Consequently, solving (67) yields $\lambda^{*}=0$ when $\varsigma=0$ and $\lambda^{*}=1$ when $\xi=0$. This completes the proof of Corollary 1.
A.4 Proof of Corollary 2
Corollary 2. With Theorem 1 holding and $\beta_{l}=\beta_{u}=\beta>0$, the optimal weight $\lambda^{*}$ that minimizes the bound in (16) takes the form
$\lambda^{*}=\frac{\beta^{2}\varsigma}{\beta^{2}\varsigma+\beta\xi}.$ (68)
Proof.
Note when $\beta_{l}=\beta_{u}=\beta$ that the right hand side of (16) reduces to
$\left(\frac{1-\lambda}{1-\lambda+\lambda/\beta}\right)^{2}\varsigma+\left(\frac{\lambda}{(1-\lambda)\beta+\lambda}\right)^{2}\xi.$ (69)
i.e.,
$\left(\frac{(1-\lambda)\beta}{(1-\lambda)\beta+\lambda}\right)^{2}\varsigma+\left(\frac{\lambda}{(1-\lambda)\beta+\lambda}\right)^{2}\xi.$ (70)
Taking the derivative of the previous bound with respect to $\lambda$ yields
Derivative $=2\varsigma\left(\frac{(1-\lambda)\beta}{(1-\lambda)\beta+\lambda}\right)\frac{-\beta\left((1-\lambda)\beta+\lambda\right)-(1-\lambda)\beta(1-\beta)}{\left((1-\lambda)\beta+\lambda\right)^{2}}$
$\quad+2\xi\left(\frac{\lambda}{(1-\lambda)\beta+\lambda}\right)\frac{(1-\lambda)\beta+\lambda-\lambda(1-\beta)}{\left((1-\lambda)\beta+\lambda\right)^{2}}.$ (71)
By combining and canceling terms the previous derivative reduces to
Derivative $=2\varsigma\frac{(\lambda-1)\beta^{2}}{\left((1-\lambda)\beta+\lambda\right)^{3}}+2\xi\frac{\lambda\beta}{\left((1-\lambda)\beta+\lambda\right)^{3}}.$ (72)
Notice that in the previous equation $(1-\lambda)\beta+\lambda>0$ due to that $\beta>0$ and $\lambda\in[0,1]$. We further solve the optimal weight $\lambda^{*}$ by letting the above derivative be zero, i.e.,
$\lambda^{*}=\frac{\beta^{2}\varsigma}{\beta^{2}\varsigma+\beta\xi}.$ (73)
This completes the proof of Corollary 2.
A.5 Proof of Theorem 2
Theorem 2 (Worst-Case Performance Bound). Under the hypotheses of Theorem 1, the following bound holds with probability at least $1-\delta$
$\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})-\mathcal{E}_{\mathcal{D}}(Q^{k+1})$
$\leq\left(\frac{1-\lambda}{1-\lambda+\lambda/\beta_{u}}\right)^{2}\varsigma+\left(\frac{\lambda}{(1-\lambda)\beta_{l}+\lambda}\right)^{2}\xi$
$\quad+\sqrt{\frac{1}{2}\log(\frac{1}{\delta})}\frac{|\mathcal{S}||\mathcal{A}|}{\sqrt{N}}\left(\frac{8\gamma B^{2}}{(1-\gamma)^{2}}\frac{\beta_{u}(1-\lambda)^{2}}{(1-\lambda)\beta_{l}^{2}+\lambda\beta_{l}}+\frac{4\gamma B}{1-\gamma}\frac{\lambda(1-\lambda)\sqrt{\xi}}{(1-\lambda)\beta_{l}^{2}+\lambda\beta_{l}}\right).$ (74)
Proof.
Note that
$\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})-\mathcal{E}_{\mathcal{D}}(Q^{k+1})$ (75)
$=\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})-\underset{\begin{subarray}{c}N(s,a),\,\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\!\left[\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})\right]+\underset{\begin{subarray}{c}N(s,a),\,\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\!\left[\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})\right]-\mathcal{E}_{\mathcal{D}}(Q^{k+1})$ (76)
$\leq\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})-\underset{\begin{subarray}{c}N(s,a),\,\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\!\left[\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})\right]+\left(\frac{1-\lambda}{1-\lambda+\lambda/\beta_{u}}\right)^{2}\varsigma+\left(\frac{\lambda}{(1-\lambda)\beta_{l}+\lambda}\right)^{2}\xi,$ (77)
where the previous inequality follows from Theorem 1. Thus, we are left to bound $\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})-\underset{\begin{subarray}{c}N(s,a),\,\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\left[\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})\right]$. To proceed, we rely on the following technical lemma.
Lemma 1 (McDiarmid Inequality).
Let $\tau_{1},\cdots,\tau_{n}$ be independent random variables taking on values in a set $H$ and let $c_{1},\cdots,c_{n}$ be positive real constants. If $\varphi:H^{n}\to\mathbb{R}$ satisfies
$\sup_{\tau_{1},\cdots,\tau_{n},\tau_{i}^{\prime}\in H}|\varphi(\tau_{1},\cdots,\tau_{i},\cdots,\tau_{n})-\varphi(\tau_{1},\cdots,\tau_{i}^{\prime},\cdots,\tau_{n})|\leq c_{i},$ (78)
for $1\leq i\leq n$, then it holds that
$P\left(\varphi(\tau_{1},\cdots,\tau_{n})-\mathbb{E}\left[\varphi(\tau_{1},\cdots,\tau_{n})\right]\geq\epsilon\right)\leq\exp\left(\frac{-2\epsilon^{2}}{\sum_{i=1}^{n}c_{i}^{2}}\right).$ (79)
To obtain the similar generalization bound akin to the above, we aim to compute the bound of $\left|\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})-\mathcal{E}_{\mathcal{D}}(\hat{Q}_{\lambda}^{k+1})\right|$, where $Q_{\lambda}^{k+1}$ and $\hat{Q}_{\lambda}^{k+1}$ (both see (13)) differ in a single sample of $\hat{s}^{\prime}$ only. More specifically, they take the sequences of random samples $(\hat{s}^{\prime}_{1},\cdots,\hat{s}^{\prime}_{i},\cdots,\hat{s}^{\prime}_{N})$ and $(\hat{s}^{\prime\prime}_{1},\cdots,\hat{s}^{\prime\prime}_{i},\cdots,\hat{s}^{\prime\prime}_{N})$ respectively, where $\hat{s}^{\prime}_{j}=\hat{s}^{\prime\prime}_{j}$ for any $j\in\{1,2,\cdots,N\}$ and $j\neq i$. Note that
$\left|\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})-\mathcal{E}_{\mathcal{D}}(\hat{Q}_{\lambda}^{k+1})\right|$
=$\stackrel{{\scriptstyle(a)}}{{=}}\left|\underset{(s,a,s^{\prime})\sim\mathcal{D}}{\mathbb{E}}\left[\left(Q_{\lambda}^{k+1}(s,a)-\hat{\mathcal{B}}_{s^{\prime}}Q^{k}(s,a)\right)^{2}-\left(\hat{Q}_{\lambda}^{k+1}(s,a)-\hat{\mathcal{B}}_{s^{\prime}}Q^{k}(s,a)\right)^{2}\right]\right|$ (80)
=$\stackrel{{\scriptstyle(b)}}{{=}}\left|\underset{(s,a,s^{\prime})\sim\mathcal{D}}{\mathbb{E}}\left[(Q_{\lambda}^{k+1}(s,a))^{2}-(\hat{Q}_{\lambda}^{k+1}(s,a))^{2}+2\hat{\mathcal{B}}_{s^{\prime}}Q^{k}(s,a)\left(\hat{Q}_{\lambda}^{k+1}(s,a)-Q_{\lambda}^{k+1}(s,a)\right)\right]\right|,$ (81)
where $(a)$ follows from the definition in (4) and (5), $(b)$ follows from expanding the square and canceling the terms.
Notice that ^$\hat{\mathcal{B}}_{s^{\prime}}Q^{k}(s,a)$ is the only term in the previous expression that depends on $s^{\prime}$. To proceed, we condition on $(s,a)$ in the previous equation, i.e.,
$\left|\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})-\mathcal{E}_{\mathcal{D}}(\hat{Q}_{\lambda}^{k+1})\right|$
=$\stackrel{{\scriptstyle(a)}}{{=}}\left|\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[(Q_{\lambda}^{k+1}(s,a))^{2}-(\hat{Q}_{\lambda}^{k+1}(s,a))^{2}+2\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)\left(\hat{Q}_{\lambda}^{k+1}(s,a)-Q_{\lambda}^{k+1}(s,a)\right)\right]\right|$ (82)
=$\stackrel{{\scriptstyle(b)}}{{=}}\left|\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[\left(Q_{\lambda}^{k+1}(s,a)-\hat{Q}_{\lambda}^{k+1}(s,a)\right)\left(Q_{\lambda}^{k+1}(s,a)+\hat{Q}_{\lambda}^{k+1}(s,a)-2\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)\right)\right]\right|,$ (83)
$\stackrel{{\scriptstyle(c)}}{{\leq}}\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[\underbrace{\left|\left(Q_{\lambda}^{k+1}(s,a)-\hat{Q}_{\lambda}^{k+1}(s,a)\right)\right|}_{G_{1}}\cdot\underbrace{\left|\left(Q_{\lambda}^{k+1}(s,a)+\hat{Q}_{\lambda}^{k+1}(s,a)-2\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)\right)\right|}_{G_{2}}\right],$ (84)
where $(a)$ follows by definition $\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)=\mathbb{E}_{s^{\prime}\sim P_{\mathcal{D}}(s^{\prime}|s,a)}\left[\hat{\mathcal{B}}_{s^{\prime}}Q^{k}(s,a)\right]$ (see (2) or (3)), $(b)$ follows by combining the terms, and $(c)$ follows by Jensen’s inequality.
We next work on $G_{1}$ and $G_{2}$ individually. Substituting the expressions of $Q_{\lambda}^{k+1}(s,a)$ and $\hat{Q}_{\lambda}^{k+1}(s,a)$ from (13) into $G_{1}$ and $G_{2}$ yields
$G_{1}$ $=\left|\frac{(1-\lambda)/N\left(\hat{\mathcal{B}}_{\hat{s}_{i}^{\prime}}Q^{k}(s,a)-\hat{\mathcal{B}}_{\hat{s}_{i}^{\prime\prime}}Q^{k}(s,a)\right)}{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)+\lambda P_{\mathcal{D}^{\prime}}(s,a)}\right|.$ (85)
$G_{2}\!=\!\left|\frac{\frac{1-\lambda}{N}\sum_{j=1}^{N(s,a)}\!\left(\hat{\mathcal{B}}_{\hat{s}_{j}^{\prime}}Q^{k}(s,a)+\hat{\mathcal{B}}_{\hat{s}_{j}^{\prime\prime}}Q^{k}(s,a)\right)+2\lambda P_{\mathcal{D}^{\prime}}(s,a)\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)}{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)+\lambda P_{\mathcal{D}^{\prime}}(s,a)}-2\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)\right|.$ (86)
By combining and canceling the terms $G_{2}$ can be further simplified as
$G_{2}$ $=\Bigg{|}\frac{\frac{1-\lambda}{N}\sum_{j=1}^{N(s,a)}\left(\hat{\mathcal{B}}_{\hat{s}_{j}^{\prime}}Q^{k}(s,a)+\hat{\mathcal{B}}_{\hat{s}_{j}^{\prime\prime}}Q^{k}(s,a)\right)-2(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)}{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)+\lambda P_{\mathcal{D}^{\prime}}(s,a)}$
$\quad\quad+\frac{2\lambda P_{\mathcal{D}^{\prime}}(s,a)\left(\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)-\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)\right)}{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)+\lambda P_{\mathcal{D}^{\prime}}(s,a)}\Bigg{|}.$ (87)
Note that Assumption 1 and (11) combining with (2) or (3) implies
|^$\left|\hat{\mathcal{B}}_{\hat{s}_{i}^{\prime}}Q^{k}(s,a)-\hat{\mathcal{B}}_{\hat{s}_{i}^{\prime\prime}}Q^{k}(s,a)\right|\leq\gamma(\frac{B}{1-\gamma}+\frac{B}{1-\gamma})=\frac{2\gamma B}{1-\gamma},\,\forall i$ (88)
|^$\left|\hat{\mathcal{B}}_{\hat{s}_{j}^{\prime}}Q^{k}(s,a)\right|\leq B+\gamma\frac{B}{1-\gamma}=\frac{B}{1-\gamma},\,\forall j$ (89)
$\left|\sum_{j=1}^{N(s,a)}\hat{\mathcal{B}}_{\hat{s}_{j}^{\prime}}Q^{k}(s,a)\right|\leq\sum_{j=1}^{N(s,a)}\left|\hat{\mathcal{B}}_{\hat{s}_{j}^{\prime}}Q^{k}(s,a)\right|\leq N(s,a)\frac{B}{1-\gamma}.$ (90)
Applying the above bounds to $G_{1}$ and $G_{2}$ yields
$G_{1}\leq\left|\frac{(1-\lambda)/N}{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)+\lambda P_{\mathcal{D}^{\prime}}(s,a)}\frac{2\gamma B}{1-\gamma}\right|.$ (91)
$G_{2}$ $\leq\Bigg{|}\frac{2\frac{1-\lambda}{N}N(s,a)\frac{B}{1-\gamma}-2(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)}{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)+\lambda P_{\mathcal{D}^{\prime}}(s,a)}$
$\quad\quad+\frac{2\lambda P_{\mathcal{D}^{\prime}}(s,a)\left(\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)-\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)\right)}{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)+\lambda P_{\mathcal{D}^{\prime}}(s,a)}\Bigg{|}$ (92)
$\leq\Bigg{|}\frac{2(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)(\frac{B}{1-\gamma}+\frac{B}{1-\gamma})}{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)+\lambda P_{\mathcal{D}^{\prime}}(s,a)}+\frac{2\lambda P_{\mathcal{D}^{\prime}}(s,a)\left(\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)-\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)\right)}{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)+\lambda P_{\mathcal{D}^{\prime}}(s,a)}\Bigg{|}$ (93)
$=\left|\frac{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)(4B/(1-\gamma))+2\lambda P_{\mathcal{D}^{\prime}}(s,a)\left(\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)-\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)\right)}{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)+\lambda P_{\mathcal{D}^{\prime}}(s,a)}\right|.$ (94)
Substituting $G_{1}$ and $G_{2}$ back into (84) yields
$\left|\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})-\mathcal{E}_{\mathcal{D}}(\hat{Q}_{\lambda}^{k+1})\right|$
$\leq\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\Bigg{[}\left|\frac{(1-\lambda)/N}{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)+\lambda P_{\mathcal{D}^{\prime}}(s,a)}\frac{2\gamma B}{1-\gamma}\right|$
$\quad\cdot\left|\frac{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)(4B/(1-\gamma))+2\lambda P_{\mathcal{D}^{\prime}}(s,a)\left|\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)-\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)\right|}{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)+\lambda P_{\mathcal{D}^{\prime}}(s,a)}\right|\Bigg{]}.$ (95)
We rewrite the previous inequality using the definition of the expectation
$\left|\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})-\mathcal{E}_{\mathcal{D}}(\hat{Q}_{\lambda}^{k+1})\right|$
$\leq\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}P_{\mathcal{D}}(s,a)\left(\frac{(1-\lambda)/N}{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)+\lambda P_{\mathcal{D}^{\prime}}(s,a)}\frac{2\gamma B}{1-\gamma}\right)$
$\quad\cdot\left|\left(\frac{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)(4B/(1-\gamma))+2\lambda P_{\mathcal{D}^{\prime}}(s,a)\left|\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)-\mathcal{B}_{\mathcal{D}}Q^{k}(s,a))\right|}{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)+\lambda P_{\mathcal{D}^{\prime}}(s,a)}\right)\right|$ (96)
$=\frac{1}{N}\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}\left(\frac{P_{\mathcal{D}}(s,a)}{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)+\lambda P_{\mathcal{D}^{\prime}}(s,a)}\frac{2(1-\lambda)\gamma B}{1-\gamma}\right)$
$\quad\cdot\left|\left(\frac{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)(4B/(1-\gamma))+2\lambda P_{\mathcal{D}^{\prime}}(s,a)\left|\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)-\mathcal{B}_{\mathcal{D}}Q^{k}(s,a))\right|}{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)+\lambda P_{\mathcal{D}^{\prime}}(s,a)}\right)\right|.$ (97)
Dividing both the numerator and denominator by $P_{\mathcal{D}}(s,a)$ in the first term and by $P_{\mathcal{D}^{\prime}}(s,a)$ in the second term of the right hand side of the previous expression, and re-ordering terms yields
$\left|\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})-\mathcal{E}_{\mathcal{D}}(\hat{Q}_{\lambda}^{k+1})\right|$
$\leq\frac{1}{N}\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}\left(\frac{1}{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)/P_{\mathcal{D}}(s,a)+\lambda P_{\mathcal{D}^{\prime}}(s,a)/P_{\mathcal{D}}(s,a)}\frac{2(1-\lambda)\gamma B}{1-\gamma}\right)$
$\cdot\left|\left(\frac{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)/P_{\mathcal{D}^{\prime}}(s,a)(4B/(1-\gamma))+2\lambda\left|\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)-\mathcal{B}_{\mathcal{D}}Q^{k}(s,a))\right|}{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)/P_{\mathcal{D}^{\prime}}(s,a)+\lambda}\right)\right|.$ (98)
By using Assumption 2 the previous inequality reduces to
$\left|\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})-\mathcal{E}_{\mathcal{D}}(\hat{Q}_{\lambda}^{k+1})\right|$
$\leq\frac{1}{N}\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}\left(\frac{1}{(1-\lambda)\beta_{l}+\lambda\beta_{l}}\frac{2(1-\lambda)\gamma B}{1-\gamma}\right)$
$\cdot\left|\left(\frac{(1-\lambda)\beta_{u}(4B/(1-\gamma))+2\lambda\left|\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a)-\mathcal{B}_{\mathcal{D}}Q^{k}(s,a))\right|}{(1-\lambda)\beta_{l}+\lambda}\right)\right|$ (99)
$\leq\frac{1}{N}\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}\left(\frac{1}{\beta_{l}}\frac{2(1-\lambda)\gamma B}{1-\gamma}\right)$
$\cdot\left|\left(\frac{(1-\lambda)\beta_{u}(4B/(1-\gamma))+2\lambda\underset{(s,a)\in\mathcal{S}\times\mathcal{A}}{\max}\left|\mathcal{B}_{\mathcal{D}}Q^{k}(s,a)-\mathcal{B}_{\mathcal{D^{\prime}}}Q^{k}(s,a))\right|}{(1-\lambda)\beta_{l}+\lambda}\right)\right|.$ (100)
By the definition of $\xi$ as in (14), the above expression is equivalent to
$\left|\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})-\mathcal{E}_{\mathcal{D}}(\hat{Q}_{\lambda}^{k+1})\right|$
$\leq\frac{1}{N}\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}\left(\frac{1}{\beta_{l}}\frac{2(1-\lambda)\gamma B}{1-\gamma}\right)\cdot\left|\left(\frac{(1-\lambda)\beta_{u}(4B/(1-\gamma))+2\lambda\sqrt{\xi}}{(1-\lambda)\beta_{l}+\lambda}\right)\right|$ (101)
$\leq\frac{|\mathcal{S}||\mathcal{A}|}{N}\left(\frac{8\gamma B^{2}}{(1-\gamma)^{2}}\frac{\beta_{u}(1-\lambda)^{2}}{(1-\lambda)\beta_{l}^{2}+\lambda\beta_{l}}+\frac{4\gamma B}{1-\gamma}\frac{\lambda(1-\lambda)\sqrt{\xi}}{(1-\lambda)\beta_{l}^{2}+\lambda\beta_{l}}\right)$ (102)
$=c,$ (103)
where $c$ is the bound as shown in Lemma 1. Lemma 1 implies that
$P\left(\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})-\underset{\begin{subarray}{c}N(s,a),\,\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\left[\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})\right]\geq\epsilon\right)\leq\exp\frac{-2\epsilon^{2}}{\sum_{i=1}^{N}c^{2}}$ (104)
i.e.,
$P\left(\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})-\underset{\begin{subarray}{c}N(s,a),\,\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\left[\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})\right]<\epsilon\right)\geq 1-\exp\frac{-2\epsilon^{2}}{Nc^{2}}.$ (105)
Let the right hand side of the previous expression to be $1-\delta$. Then,
$\epsilon$ $=\sqrt{\frac{1}{2}\log(\frac{1}{\delta})N}c$ (106)
$=\sqrt{\frac{1}{2}\log(\frac{1}{\delta})}\frac{|\mathcal{S}||\mathcal{A}|}{\sqrt{N}}\left(\frac{8\gamma B^{2}}{(1-\gamma)^{2}}\frac{\beta_{u}(1-\lambda)^{2}}{(1-\lambda)\beta_{l}^{2}+\lambda\beta_{l}}+\frac{4\gamma B}{1-\gamma}\frac{\lambda(1-\lambda)\sqrt{\xi}}{(1-\lambda)\beta_{l}^{2}+\lambda\beta_{l}}\right).$ (107)
Therefore, the following bound holds with probability at least $1-\delta$
$\mathcal{E}_{\mathcal{D}}(Q_{\lambda}^{k+1})-\mathcal{E}_{\mathcal{D}}(Q^{k+1})$
$\leq\left(\frac{1-\lambda}{1-\lambda+\lambda/\beta_{u}}\right)^{2}\varsigma+\left(\frac{\lambda}{(1-\lambda)\beta_{l}+\lambda}\right)^{2}\xi$
$\quad+\sqrt{\frac{1}{2}\log(\frac{1}{\delta})}\frac{|\mathcal{S}||\mathcal{A}|}{\sqrt{N}}\left(\frac{8\gamma B^{2}}{(1-\gamma)^{2}}\frac{\beta_{u}(1-\lambda)^{2}}{(1-\lambda)\beta_{l}^{2}+\lambda\beta_{l}}+\frac{4\gamma B}{1-\gamma}\frac{\lambda(1-\lambda)\sqrt{\xi}}{(1-\lambda)\beta_{l}^{2}+\lambda\beta_{l}}\right).$ (108)
which completes the proof of Theorem 2.
A.6 Proof of Theorem 3
To proceed, we rely on the following technical lemma.
Lemma 2.
For all iterations $(k=0,1,2,\cdots)$, recall $\xi(Q^{k})$ with respect to $Q^{k}$ as in (14), and let us define $\xi_{\text{max}}=\sup_{k\in\mathbb{N}}\,\xi(Q^{k})$ and
$\sigma_{\text{max}}=\sup_{k\in\mathbb{N}}\,\underset{\begin{subarray}{c}(s,a)\in\mathcal{S}\times\mathcal{A}\end{subarray}}{\max}\left[\underset{\hat{\mathcal{D}}}{\max}\left[||\left(\frac{1}{N(s,a)}\sum_{j=1}^{N(s,a)}\hat{\mathcal{B}}_{s_{j}^{\prime}}Q_{\lambda}^{k}(s,a)-\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a)\right)||\right]\right].$ (109)
Then, it holds that
$\underset{\begin{subarray}{c}N(s,a),\,\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\left[\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[||Q_{\lambda}^{k+1}(s,a)-\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a)||_{\infty}\right]\right]$
$\leq\frac{1-\lambda}{1-\lambda+\lambda/\beta_{u}}\sigma_{\text{max}}+\frac{\lambda}{(1-\lambda)\beta_{l}+\lambda}\sqrt{\xi_{\text{max}}},~{}\forall k\in\mathbb{N}.$ (110)
Proof.
For all iterations $k=0,1,2,\cdots$,
$\underset{\begin{subarray}{c}N(s,a),\,\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\left[\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[||Q_{\lambda}^{k+1}(s,a)-\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a)||_{\infty}\right]\right]$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[\underset{\begin{subarray}{c}N(s,a),\,\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\left[||Q_{\lambda}^{k+1}(s,a)-\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a)||\right]\right]$ (111)
=$\stackrel{{\scriptstyle(b)}}{{=}}\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\!\!\left[\underset{\begin{subarray}{c}N(s,a)\\ \hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\!\left[||\frac{\frac{1-\lambda}{N}\sum_{j=1}^{N(s,a)}\!\hat{\mathcal{B}}_{s_{j}^{\prime}}Q_{\lambda}^{k}(s,a)\!+\!\lambda P_{\mathcal{D}^{\prime}}(s,a)\mathcal{B}_{\mathcal{D^{\prime}}}Q_{\lambda}^{k}(s,a)}{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)+\lambda P_{\mathcal{D}^{\prime}}(s,a)}\!-\!\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a)||\right]\!\right]$ (112)
=$\stackrel{{\scriptstyle(c)}}{{=}}\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\Bigg{[}\underset{\begin{subarray}{c}N(s,a)\\ \hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\Bigg{[}||\frac{\frac{1-\lambda}{N}\sum_{j=1}^{N(s,a)}\hat{\mathcal{B}}_{s_{j}^{\prime}}Q_{\lambda}^{k}(s,a)-(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a))}{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)+\lambda P_{\mathcal{D}^{\prime}}(s,a)}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad+\frac{\lambda P_{\mathcal{D}^{\prime}}(s,a)(\mathcal{B}_{\mathcal{D^{\prime}}}Q_{\lambda}^{k}(s,a)-\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a))}{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)+\lambda P_{\mathcal{D}^{\prime}}(s,a)}||\Bigg{]}\Bigg{]}$ (113)
where $(a)$ follows from swapping the two expectations and the fact that $||\cdot||_{\infty}\leq||\cdot||$, $(b)$ follows by definition of $Q_{\lambda}^{k+1}(s,a)$ (see (13)) that operates over $Q_{\lambda}^{k}(s,a)$, $(c)$ follows by combining the terms.
By using the triangle inequality the previous expression reduces to
$\underset{\begin{subarray}{c}N(s,a),\,\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\left[\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[||Q_{\lambda}^{k+1}(s,a)-\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a)||_{\infty}\right]\right]$
$\leq\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[\underset{\begin{subarray}{c}N(s,a)\\ \hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\left[||\frac{\frac{1-\lambda}{N}\sum_{j=1}^{N(s,a)}\hat{\mathcal{B}}_{s_{j}^{\prime}}Q_{\lambda}^{k}(s,a)-(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a))}{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)+\lambda P_{\mathcal{D}^{\prime}}(s,a)}||\right]\right]$
$\quad+\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[\underset{\begin{subarray}{c}N(s,a)\\ \hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\left[||\frac{\lambda P_{\mathcal{D}^{\prime}}(s,a)(\mathcal{B}_{\mathcal{D^{\prime}}}Q_{\lambda}^{k}(s,a)-\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a))}{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)+\lambda P_{\mathcal{D}^{\prime}}(s,a)}||\right]\right].$ (114)
By the definition $P_{\hat{\mathcal{D}}}(s,a)=N(s,a)/N$, the previous inequality is equivalent to
$\underset{\begin{subarray}{c}N(s,a),\,\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\left[\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[||Q_{\lambda}^{k+1}(s,a)-\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a)||_{\infty}\right]\right]$
$\leq\!\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\!\!\left[\underset{\begin{subarray}{c}N(s,a)\\ \hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\!\!\left[\!||\frac{(1-\lambda)\frac{P_{\hat{\mathcal{D}}}(s,a)}{N(s,a)}\!\sum_{j=1}^{N(s,a)}\!\hat{\mathcal{B}}_{s_{j}^{\prime}}Q_{\lambda}^{k}(s,a)\!-\!(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a))}{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)+\lambda P_{\mathcal{D}^{\prime}}(s,a)}||\!\right]\!\right]$
$\quad+\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[\underset{\begin{subarray}{c}N(s,a)\\ \hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\left[||\frac{\lambda P_{\mathcal{D}^{\prime}}(s,a)(\mathcal{B}_{\mathcal{D^{\prime}}}Q_{\lambda}^{k}(s,a)-\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a))}{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)+\lambda P_{\mathcal{D}^{\prime}}(s,a)}||\right]\right].$ (115)
Dividing both the numerator and denominator by $P_{\hat{\mathcal{D}}}(s,a)$ in the first term and by $P_{\mathcal{D}^{\prime}}(s,a)$ in the second term of the right hand side of the previous inequality, and re-ordering terms yields
$\underset{\begin{subarray}{c}N(s,a),\,\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\left[\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[||Q_{\lambda}^{k+1}(s,a)-\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a)||_{\infty}\right]\right]$
$\leq\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[\underset{\begin{subarray}{c}N(s,a)\\ \hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\left[||\frac{(1-\lambda)\left(\frac{1}{N(s,a)}\sum_{j=1}^{N(s,a)}\hat{\mathcal{B}}_{s_{j}^{\prime}}Q_{\lambda}^{k}(s,a)-\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a)\right)}{1-\lambda+\lambda P_{\mathcal{D}^{\prime}}(s,a)/P_{\hat{\mathcal{D}}}(s,a)}||\right]\right]$
$\quad+\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[\underset{\begin{subarray}{c}N(s,a)\\ \hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\left[||\frac{\lambda(\mathcal{B}_{\mathcal{D^{\prime}}}Q_{\lambda}^{k}(s,a)-\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a))}{(1-\lambda)P_{\hat{\mathcal{D}}}(s,a)/P_{\mathcal{D}^{\prime}}(s,a)+\lambda}||\right]\right].$ (116)
Assumption 2 implies that both $P_{\hat{\mathcal{D}}}(s,a)/P_{\mathcal{D}^{\prime}}(s,a)$ and $P_{\mathcal{D}^{\prime}}(s,a)/P_{\hat{\mathcal{D}}}(s,a)$ are bounded, i.e.,
$\beta_{l}\leq\frac{P_{\hat{\mathcal{D}}}(s,a)}{P_{\mathcal{D}^{\prime}}(s,a)}\leq\beta_{u},\quad\frac{1}{\beta_{u}}\leq\frac{P_{\mathcal{D}^{\prime}}(s,a)}{P_{\hat{\mathcal{D}}}(s,a)}\leq\frac{1}{\beta_{l}}.$ (117)
Applying the previous bounds to (116) yields
$\underset{\begin{subarray}{c}N(s,a),\,\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\left[\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[||Q_{\lambda}^{k+1}(s,a)-\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a)||_{\infty}\right]\right]$
$\leq\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[\underset{\begin{subarray}{c}N(s,a)\\ \hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\left[||\frac{(1-\lambda)\left(\frac{1}{N(s,a)}\sum_{j=1}^{N(s,a)}\hat{\mathcal{B}}_{s_{j}^{\prime}}Q_{\lambda}^{k}(s,a)-\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a)\right)}{1-\lambda+\lambda/\beta_{u}}||\right]\right]$
$\quad+\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[\underset{\begin{subarray}{c}N(s,a)\\ \hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\left[||\frac{\lambda(\mathcal{B}_{\mathcal{D^{\prime}}}Q_{\lambda}^{k}(s,a)-\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a))}{(1-\lambda)\beta_{l}+\lambda}||\right]\right].$ (118)
Extracting the constant terms in the above expression outside expectations and using the fact that $\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a)$ and $\mathcal{B}_{\mathcal{D^{\prime}}}Q_{\lambda}^{k}(s,a)$ do not depend on $\hat{s}^{\prime}$ (at the step $k+1$) yields
$\underset{\begin{subarray}{c}N(s,a),\,\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\left[\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[||Q_{\lambda}^{k+1}(s,a)-\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a)||_{\infty}\right]\right]$
$\leq\frac{1-\lambda}{1-\lambda+\lambda/\beta_{u}}\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[\underset{\begin{subarray}{c}N(s,a)\\ \hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\left[||\left(\frac{1}{N(s,a)}\sum_{j=1}^{N(s,a)}\hat{\mathcal{B}}_{s_{j}^{\prime}}Q_{\lambda}^{k}(s,a)-\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a)\right)||\right]\right]$
$\quad+\frac{\lambda}{(1-\lambda)\beta_{l}+\lambda}\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[||(\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a)-\mathcal{B}_{\mathcal{D^{\prime}}}Q_{\lambda}^{k}(s,a))||\right]$ (119)
$\leq\frac{1-\lambda}{1-\lambda+\lambda/\beta_{u}}\underset{\begin{subarray}{c}(s,a)\in\mathcal{S}\times\mathcal{A}\end{subarray}}{\max}\left[\underset{\hat{\mathcal{D}}}{\max}\left[||\left(\frac{1}{N(s,a)}\sum_{j=1}^{N(s,a)}\hat{\mathcal{B}}_{s_{j}^{\prime}}Q_{\lambda}^{k}(s,a)-\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a)\right)||\right]\right]$
$\quad+\frac{\lambda}{(1-\lambda)\beta_{l}+\lambda}\underset{\begin{subarray}{c}(s,a)\in\mathcal{S}\times\mathcal{A}\end{subarray}}{\max}\left[||(\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a)-\mathcal{B}_{\mathcal{D^{\prime}}}Q_{\lambda}^{k}(s,a))||\right]$ (120)
By definition of $\sigma_{\text{max}}$ and $\xi_{\text{max}}$ we obtain
$\underset{\begin{subarray}{c}N(s,a),\,\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\left[\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[||Q_{\lambda}^{k+1}(s,a)-\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a)||_{\infty}\right]\right]$
$\leq\frac{1-\lambda}{1-\lambda+\lambda/\beta_{u}}\sigma_{\text{max}}+\frac{\lambda}{(1-\lambda)\beta_{l}+\lambda}\sqrt{\xi_{\text{max}}},~{}\forall k\in\mathbb{N}.$ (121)
This completes the proof of Lemma 2. ∎
Having introduced Lemma 2, we are in the stage of proving Theorem 3.
Theorem 3 (Convergence). Under the hypotheses of Theorem 1, it holds that
$\underset{\begin{subarray}{c}N(s,a),\,\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\left[\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\!\!\left[||Q_{\lambda}^{k+1}(s,a)\!-\!Q^{*}(s,a)||_{\infty}\right]\right]\leq$
$\gamma^{k+1}\!\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[||Q^{0}(s,a)-Q^{*}(s,a)||_{\infty}\right]\!+\!\frac{1-\gamma^{k+1}}{1-\gamma}\!\!\left(\frac{1-\lambda}{1-\lambda+\lambda/\beta_{u}}\sigma_{\text{max}}\!+\!\frac{\lambda}{(1-\lambda)\beta_{l}+\lambda}\sqrt{\xi_{\text{max}}}\right).$ (122)
Proof.
From the triangle inequality we obtain
$\underset{\begin{subarray}{c}N(s,a),\,\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\left[\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[||Q_{\lambda}^{k+1}(s,a)-Q^{*}(s,a)||_{\infty}\right]\right]$ (123)
$\leq\underset{\begin{subarray}{c}N(s,a),\,\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\!\left[\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[||Q_{\lambda}^{k+1}(s,a)-\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a)||_{\infty}\!+\!||\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a)-Q^{*}(s,a)||_{\infty}\right]\right].$ (124)
Given that $Q^{*}$ is the fixed point of the Bellman optimality operator (2) (i.e., $\mathcal{B}_{\mathcal{D}}Q^{*}(s,a)=Q^{*}(s,a)$), we obtain
$\underset{\begin{subarray}{c}N(s,a),\,\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\left[\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[||Q_{\lambda}^{k+1}(s,a)-Q^{*}(s,a)||_{\infty}\right]\right]$
$\leq\!\underset{\begin{subarray}{c}N(s,a),\,\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\!\!\left[\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[||Q_{\lambda}^{k+1}(s,a)\!-\!\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a)||_{\infty}\!+\!||\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a)\!-\!\mathcal{B}_{\mathcal{D}}Q^{*}(s,a)||_{\infty}\right]\!\right].$ (125)
Applying the contraction property  (Sutton & Barto, 2018) of the Bellman optimality operator to the previous inequality and re-arranging the terms yields
$\underset{\begin{subarray}{c}N(s,a),\,\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\left[\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[||Q_{\lambda}^{k+1}(s,a)-Q^{*}(s,a)||_{\infty}\right]\right]$
$\leq\underset{\begin{subarray}{c}N(s,a),\,\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\left[\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[||Q_{\lambda}^{k+1}(s,a)-\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a)||_{\infty}+\gamma||Q_{\lambda}^{k}(s,a)-Q^{*}(s,a)||_{\infty}\right]\right]$ (126)
$=\underset{\begin{subarray}{c}N(s,a),\,\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\left[\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[||Q_{\lambda}^{k+1}(s,a)-\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a)||_{\infty}\right]\right]$
$\quad+\gamma\underset{\begin{subarray}{c}N(s,a),\,\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\left[\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[||Q_{\lambda}^{k}(s,a)-Q^{*}(s,a)||_{\infty}\right]\right].$ (127)
Unrolling the previous inequality until $Q_{\lambda}^{0}$ yields
$\underset{\begin{subarray}{c}N(s,a),\,\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\left[\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[||Q_{\lambda}^{k+1}(s,a)-Q^{*}(s,a)||_{\infty}\right]\right]$
$\leq\underset{\begin{subarray}{c}N(s,a),\,\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\left[\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[||Q_{\lambda}^{k+1}(s,a)-\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k}(s,a)||_{\infty}\right]\right]$
$+\gamma\underset{\begin{subarray}{c}N(s,a),\,\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\left[\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[||Q_{\lambda}^{k}(s,a)-\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{k-1}(s,a)||_{\infty}\right]\right]+\cdots$
$+\gamma^{k}\underset{\begin{subarray}{c}N(s,a),\,\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\left[\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[||Q_{\lambda}^{1}(s,a)-\mathcal{B}_{\mathcal{D}}Q_{\lambda}^{0}(s,a)||_{\infty}\right]\right]$
$+\gamma^{k+1}\underset{\begin{subarray}{c}N(s,a),\,\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\left[\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[||Q_{\lambda}^{0}(s,a)-Q^{*}(s,a)||_{\infty}\right]\right].$ (128)
Since $Q^{0}$ is the initial Q-function, it holds that
$Q^{0}(s,a)=Q_{\lambda}^{0}(s,a),\,\forall(s,a)\in\mathcal{S}\times\mathcal{A}.$ (129)
Employing Lemma 2 further implies that
$\underset{\begin{subarray}{c}N(s,a),\,\hat{s}^{\prime}\sim P_{\mathcal{D}}(\hat{s}^{\prime}|s,a)\end{subarray}}{\mathbb{E}}\left[\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[||Q_{\lambda}^{k+1}(s,a)-Q^{*}(s,a)||_{\infty}\right]\right]\leq$
$\left(\frac{1-\lambda}{1-\lambda+\lambda/\beta_{u}}\sigma_{\text{max}}\!+\!\frac{\lambda}{(1-\lambda)\beta_{l}+\lambda}\sqrt{\xi_{\text{max}}}\right)\!+\!\gamma\left(\frac{1-\lambda}{1-\lambda+\lambda/\beta_{u}}\sigma_{\text{max}}\!+\!\frac{\lambda}{(1-\lambda)\beta_{l}+\lambda}\sqrt{\xi_{\text{max}}}\right)$
$+\cdots+\gamma^{k}\left(\frac{1-\lambda}{1-\lambda+\lambda/\beta_{u}}\sigma_{\text{max}}+\frac{\lambda}{(1-\lambda)\beta_{l}+\lambda}\sqrt{\xi_{\text{max}}}\right)$
$+\gamma^{k+1}\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[||Q^{0}(s,a)-Q^{*}(s,a)||_{\infty}\right]$ (130)
$=\gamma^{k+1}\underset{(s,a)\sim\mathcal{D}}{\mathbb{E}}\left[||Q^{0}(s,a)-Q^{*}(s,a)||_{\infty}\right]$
$\quad+\frac{1-\gamma^{k+1}}{1-\gamma}\left(\frac{1-\lambda}{1-\lambda+\lambda/\beta_{u}}\sigma_{\text{max}}+\frac{\lambda}{(1-\lambda)\beta_{l}+\lambda}\sqrt{\xi_{\text{max}}}\right).$ (131)
This completes the proof of Theorem 3. ∎
Appendix B Environments
We select five Procgen games (Cobbe et al., 2020) to substantiate our theoretical contributions in this work, whose details are provided below.
Description of Caveflyer (Cobbe et al., 2020).
“The player needs to traverse a complex network of caves to reach the exit. Player movement is reminiscent of the classic Atari game “Asteroids” where the ship can rotate and propel forward or backward along its current axis. The primary reward is granted upon successfully reaching the end of the level, though additional reward can be earned by destroying target objects with the ship’s lasers along the way. The level is fraught with both stationary and moving lethal obstacles, demanding precise navigation and quick reflexes.”
Figure 5: The screenshot of Caveflyer (Cobbe et al., 2020).
Description of Climber (Cobbe et al., 2020).
“The player needs to climb a series of platforms, collecting stars scattered along the path. A small reward is granted for each star collected, with a substantial reward provided for gathering all stars within a level. If every star is collected, the episode terminates. The level is also populated with lethal flying monsters, adding extra challenges to the player’s journey.”
Figure 6: The screenshot of Climber (Cobbe et al., 2020).
Description of Dodgeball (Cobbe et al., 2020).
“Inspired by the Atari game “Berzerk”, the player spawns in a room with a randomly generated configuration of walls and enemies. Contact with a wall results in an immediate game over, terminating the episode. The player moves slowly, allowing for careful navigation throughout the room. Enemies, moving slowly too, throw balls at the player. The player can retaliate by throwing balls as well, but only in the direction they are facing. Once all enemies are eliminated, the player can advance to the unlocked platform, earning a substantial level completion bonus.”
Figure 7: The screenshot of Dodgeball (Cobbe et al., 2020).
Description of Maze (Cobbe et al., 2020).
“The player, embodying a mouse, needs to navigate a maze to locate the sole piece of cheese and obtain a reward. The mazes, generated using Kruskal’s algorithm, vary in size from $3\times 3$ to $25\times 25$, with dimensions uniformly sampled across this range. To navigate the maze, the player can move up, down, left, or right.”
Figure 8: The screenshot of Maze (Cobbe et al., 2020).
Description of Miner (Cobbe et al., 2020).
“Inspired by the game “BoulderDash”, the player (robot) can dig through dirt to navigate the world. The game world is governed by gravity, where dirt supports both boulders and diamonds. Boulders and diamonds fall through free spaces and roll off each other. If either a boulder or a diamond falls on the player, the game terminates immediately. The objective is to collect all the diamonds in the level and then reach the exit. The player earns a small reward for each diamond collected and a huge reward for successful completion in the level.”
Figure 9: The screenshot of Miner (Cobbe et al., 2020).
Appendix C Experimental Details
C.1 Omitted Results in the Main Text
The experimental results that are omitted in the main text are presented in Tables 1-4.
Table 1: The performance across all $\lambda\in\{0,0.2,0.4,0.5,0.6,0.8,1.0\}$ (mean and standard deviation) with fixed $N=1000$ corresponding to different $\xi$.
Game $\xi_{1}([0,99])$ $\xi_{2}([25,124])$ $\xi_{3}([50,149])$
Caveflyer $1.23\pm 0.23$ $1.28\pm 0.26$ $\mathbf{1.37\pm 0.34}$
Climber $0.57\pm 0.10$ $0.73\pm 0.16$ $\mathbf{0.85\pm 0.21}$
Dodgeball $0.67\pm 0.17$ $0.87\pm 0.21$ $\mathbf{1.29\pm 0.42}$
Maze $3.52\pm 0.41$ $3.97\pm 0.42$ $\mathbf{4.35\pm 0.52}$
Miner $1.27\pm 0.48$ $2.05\pm 0.55$ $\mathbf{3.12\pm 0.80}$
Table 2: The performance across all $\lambda\in\{0,0.2,0.4,0.5,0.6,0.8,1.0\}$ (mean and standard deviation) with fixed $\xi_{3}$ corresponding to different $N$.
Game $N=1000$ $N=2500$ $N=4000$
Caveflyer $1.37\pm 0.34$ $1.52\pm 0.26$ $\mathbf{1.57\pm 0.19}$
Climber $0.85\pm 0.21$ $0.89\pm 0.29$ $\mathbf{0.95\pm 0.26}$
Dodgeball $1.29\pm 0.42$ $1.34\pm 0.30$ $\mathbf{1.42\pm 0.18}$
Maze $4.35\pm 0.52$ $4.75\pm 0.59$ $\mathbf{5.21\pm 0.69}$
Miner $3.12\pm 0.80$ $3.28\pm 0.59$ $\mathbf{3.33\pm 0.75}$
Table 3: The optimal weight corresponding to different $\xi$.
Game $\xi_{1}([0,99])$ $\xi_{2}([25,124])$ $\xi_{3}([50,149])$
Caveflyer $0.4$ $0.4$ $0.5$
Climber $0.6$ $0.8$ $1.0$
Dodgeball $0.5$ $0.6$ $0.8$
Maze $0.4$ $0.6$ $0.8$
Miner $0.4$ $0.6$ $0.8$
Table 4: The optimal weight corresponding to different $N$.
[…]
Caveflyer $0.5$ $0.5$ $0.2$
Climber $1.0$ $0.8$ $0.4$
Dodgeball $0.8$ $0.6$ $0.5$
Maze $0.8$ $0.4$ $0.2$
Miner $0.8$ $0.8$ $0.6$
C.2 Implementation
Our implementations as well as the datasets that have been used in this work are based on (Mediratta et al., 2023). Instead of training on a single dataset, this work trains a CQL agent (Kumar et al., 2020) on two different datasets from the source and target domains. These two domains contain distinct environment levels, e.g., $[0,99]$ for the source domain and $[100,199]$ for the target domain.
C.3 Hyperparameters
We adhere to the hyperparameters set by the original implementation in (Mediratta et al., 2023), with the sole exception of assigning distinct weights to each of the two dataset. The main hyperparameters employed in this work are summarized in Table 5.
Table 5: Experimental hyperparameters.
Hyperparameters Value
Target domain levels $[100,199]$
Source domain levels $[0,99]$, $[25,124]$, $[50,149]$
Number of target samples ($N$) $1000,2500,4000$
Number of source samples $40000$
Weight ($\lambda$) $\{0,0.2,0.4,0.5,0.6,0.8,1\}$
Number of episodes for evaluation $500$
Learning rate $0.0005$
Batch size $256$
Neural network hidden size $256$
Discount factor ($\gamma$) $0.99$
CQL conservativeness constant ($\alpha$) $4$
Gradient norm clip $0.1$
C.4 Computing Infrastructure
All numerical experiments were implemented on a workstation with an NVIDIA RTX A5000 GPU (24 G), Intel Core i9-10900X CPU @ 3.70 GHz $\times$ 20.
C.5 Code
The codes will be made available upon the publication of this work.